### v1

# secReferences

Abstract
Knowledge distillation (KD) has been widely studied in unsupervised Industrial Image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information. In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on the MVTec 3D-AD dataset demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code is available at CRD.
1 Introduction
Industrial Anomaly Detection (AD) is an important task aimed at ensuring product quality by identifying defects in products. In real-world industrial scenarios, the majority of object samples are normal, and anomalous samples are often difficult to obtain. Hence, unsupervised AD, which relies only on normal samples for training, has become one of the key research directions in industrial applications [6, 8, 35]. Although traditional RGB-based analysis methods meet the industrial anomaly detection needs to a certain extent, they often struggle to effectively identify anomalies in complex industrial environments, such as in cases of large changes in illumination, or surface bumps that are indistinguishable from color differences. Given the limitations of RGB anomaly detection, data from 3D sensors, including depth map and point cloud, have been widely applied in AD in recent years, which results in growing research interest in unsupervised multimodal AD.
Figure 1: Comparison of our proposed Crossmodal Reverse Distillation (CRD) with other multimodal AD methods based on feature learning. The upper part contains an overview of the two previous categories of methods, and the lower part shows the overall framework of CRD.
Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [18] and M3DM [28]. These methods construct large memory banks by processing and storing the multimodal features of normal samples from the training set. During inference, the extracted features are compared with the normal features in the memory bank to determine if the samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [24, 15, 12]. The feature learning methods typically leverage pretrained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks. Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks. In other words, the abnormality representation ability is not transferred. During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
Currently, feature learning-based methods are divided into two categories as in Fig. 1: (1) Methods like AST [24] and MMRD [15] (Fig. 1 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples. However, these methods may smooth out abnormal features during fusion, reducing teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous. (2) Crossmodal Feature Mapping [12] (Fig. 1 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. But inevitably, misalignment between multimodal features results in underfitting and cause false positives in the output anomaly maps.
To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one. Each modality has its own distillation target, ensuring the sensitivity to anomalies within each modality. Additionally, to tackle the underfitting problem of Crossmodal Feature Mapping, we propose not to directly use the feature similarity from crossmodal mapping for anomaly map generation. Instead, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD), named Crossmodal Reverse Distillation (CRD). CRD naturally combines the strengths of both feature learning paradigms. First, we design a Multi-branch Distillation (MBD) that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two crossmodal assistants within each modality branch. One is Crossmodal Filter (CF), which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier (CA), which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of the test samples are calculated by integrating the anomaly maps of all branches. Experimental results on the representative 3D AD dataset demonstrate the effectiveness of our method in multimodal anomaly detection and localization. Our contributions are as follows:
  • We propose Crossmodal Reverse Distillation based on Multi-branch Distillation for multimodal industrial anomaly detection, which effectively detects anomalies in all modalities.
  • Two crossmodal assistants are introduced including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
  • Experimental results show that our proposed CRD achieves state-of-the-art performance on the multimodal industrial AD dataset.
2 Related Work
Unsupervised Anomaly Detection
Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [5, 1, 34]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [23, 2, 20], comparing test samples with stored normal features to detect anomalies. Recently, there has been significant progress in the research of synthetic anomaly images [19, 30, 21], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [25, 27, 24, 3, 13, 26, 16, 14, 33, 17] based on teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
Multimodal Anomaly Detection
With the release of multiple 3D industrial anomaly detection datasets [7, 9, 22], unsupervised multimodal AD has gradually become a topic of research. Some methods [11], such as BTF [18] and M3DM [28], follow PatchCore [23] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [10, 31, 32] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [24], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two lightweight neural networks to perform crossmodal feature mapping and locate anomalies based on the mapping results.
3 Preliminaries
Knowledge distillation (KD) is a widely recognized and researched paradigm for unsupervised industrial image AD, typically based on a teacher-student network framework. The teacher network is usually a pretrained model, while the student network is a trainable network identical or similar to the teacher network. During training, only normal samples are used, and the teacher network’s ability to represent normal features is transferred to the student network. This process is often realized using cosine similarity. Let $F_{T}$ represent the feature output of the teacher network, and $F_{S}$ represent the feature output of the student network. The optimization loss is generally expressed as
$\mathrm{Sim}(f_{1},f_{2})=\frac{{f_{1}}^{\mathsf{T}}f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
$\mathcal{L}=\sum_{i=1}^{l}\{1-\mathrm{Sim}(F_{T}^{i},F_{S}^{i})\}$ (2)
where $l$ represents the number of selected feature layers, typically set to 3. During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map, and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample. When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low. When the input is an anomalous sample, since the student has not learned the teacher’s abnormal feature representation ability, the feature distance $M$ is large in the anomalous regions, allowing for the detection of both the presence and location of anomalies.
Reverse distillation (RD) [13] is an advanced KD-based AD method. Unlike traditional forward distillation, RD uses an encoder-decoder architecture. The teacher network is a pre-trained encoder with a WideResNet 50 [29] backbone, while the student network is a trainable decoder that generates features matching the size of the teacher’s features. RD also includes a trainable bottleneck module, OCBE, to compress features between the teacher and student. During training, the student network learns to reconstruct the teacher’s normal features using normal samples. During inference, the teacher network, with its strong feature extraction ability, captures abnormal patterns in the anomalous input and generates anomaly features. At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network. As a result, even for abnormal samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
4 Crossmodal Reverse Distillation
Figure 2: Pipeline of our proposed method which introduces Knowledge Distillation to the multimodal industrial AD task. Our proposed method mainly consists of two parts: (1) Cross Reverse Distillation designs multiple branches to realize Reverse Distillation for multiple modalities, (2) Anomaly Scoring aggregates anomaly maps from multiple branches to achieve multimodal anomaly detection.
Figure 3: Information flow of the branch of Modality A, showing how Crossmodal Filter and Crossmodal Amplifier incorporate information from Modality B. Crossmodal Filter aids the student decoder in generating normal features by utilizing features from Modality B, while Crossmodal Amplifier strengthens the perception of anomalies in Modality A by amplifying anomaly signals from Modality B.
In the multimodal anomaly detection task, each input consists of multiple modalities of data. In this section, we primarily focus on RGB and Depth modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{RGB},\mathcal{I}^{Dep}\}$, where $\mathcal{I}^{RGB}$ and $\mathcal{I}^{Dep}$ are the RGB image and depth map respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
4.1 Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories:
  • Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem.
  • Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network.
The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
However, the effectiveness of these methods relies on a key assumption: the fused teacher features exhibit strong sensitivity to anomalies, which means they generate a significant difference between normal and anomalous regions. The issue is that in multimodal AD, it is common for one modality to be anomalous while another modality remains normal. In such cases, if the normal modality dominates the fusion, the abnormal information becomes smoothed. As a result, the fused teacher features may fail to capture the abnormal characteristics, remaining close to the normal representation. This causes the student network to fit the teacher’s features even in the anomalous regions, leading to false negatives in anomaly detection.
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 2, which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including RGB and depth branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality, eliminating the need for fusion at the input or feature exaction stages. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
For the design of each branch, we draw inspiration from RD. Both the RGB and depth branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec. 4.2). The teacher encoders in both branches share parameters and are based on WideResNet 50 pre-trained on ImageNet. The student decoders follow the design of RD, using a structure symmetric to that of the teacher networks. Crossmodal Filter replaces the OCBE module of RD, while Crossmodal Amplifier is placed after the student decoder to optimize the output features, enabling crossmodal mapping and information fusion between modalities. The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
DepthGAN [7] 53.8 37.2 58.0 60.3 43.0 53.4 64.2 60.1 44.3 57.7 53.2
DepthAE [7] 64.8 50.2 65.0 48.8 80.5 52.2 71.2 52.9 54.0 55.2 59.5
DepthVM [7] 51.3 55.1 47.7 58.1 61.7 71.6 45.0 42.1 59.8 62.3 55.5
BTF [18] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
M3DM [28] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
Shape-guided [11] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
AST [24] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
MMRD [15] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
CFM [12] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
CRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
Table 1: Image-level anomaly detection results I-AUROC (%) on MVTec 3D-AD with the best in bold. We abbreviate Crossmodal Feature Mapping as CFM, the same below.
4.2 Crossmodal Assistants
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final detection.
Crossmodal Filter
The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
Fig. 3 depicts the structure and information flow of Crossmodal Filter (discussed here with modality B assisting modality A as an example). Our proposed Crossmodal Filter module consists of two parts: Bottleneck Projection (BP) and a modified version of the OCBE module. First, BP module aligns the teacher encoder features of modality B to the teacher features of modality A, preventing interference in the decoder process of modality A during later feature fusion. Then, the aligned modality B features are concatenated with the original teacher features of modality A. The OCBE module from RD is modified to retain its feature compression capability while expanding the input feature dimensions to accommodate the fused features.
It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration. First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $8\times 8$ by convolution. Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$. During training, we minimize the distance between $F^{B}_{BP}$ and the teacher encoder features of modality A $F^{A}_{E}$, mapping modality B’s features into modality A’s feature space. During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
Crossmodal Amplifier
When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec. 4.1. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
As illustrated in Fig. 3, we design Crossmodal Amplifier. First, Inverted Bottleneck Projection (IBP) is designed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$. Then, the mapped features are then dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by
${F^{A}_{CA}}^{i}=\frac{\exp({w_{1}^{A}}^{i})*{F^{A}_{D}}^{i}+\exp({w_{2}^{A}}^{i})*{F^{B}_{IBP}}^{i}}{\exp({w_{1}^{A}}^{i})+\exp({w_{2}^{A}}^{i})}$ (3)
where $w_{1}^{A}$ and $w_{2}^{A}$ are trainable parameters, initially set to 1.0 during training.
Inverted Bottleneck Projection only learns to map modality B’s normal features to modality A’s feature space during training. It means that IBP is not able to make the correct feature mapping when modality B has anomalies in the test sample. Therefore, the fused feature ${F^{A}_{CA}}$ includes modality B’s anomalous information, which leads to a larger distance between the teacher encoder’s features $F^{A}_{E}$ and ${F^{A}_{CA}}$, allowing the anomaly in modality B to be detected in modality A’s branch.
In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes. We believe that this process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information in the modality during feature mapping.
4.3 Training Objectives
The overall optimization objective of our method is composed of three main components. Firstly, similar to the approach in RD, we optimize the features output by the student decoder using cosine similarity as
$\mathcal{L}^{RGB/Dep}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{D}^{RGB/Dep}}^{i})\}$ (4)
where $\mathcal{L}^{RGB}_{D}$ and $\mathcal{L}^{Dep}_{D}$ are the losses corresponding to RGB and depth branches respectively.
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{Dep/RGB}$ within these filters to the encoder features. The training losses of RGB and Depth branches are calculated as:
$\mathcal{L}^{RGB/Dep}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{BP}^{Dep/RGB}}^{i})\}$ (5)
Finally, Crossmodal Amplifier needs to be optimized,, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. The total loss function is
$\mathcal{L}^{RGB/Dep}_{CA}=\mathcal{L}^{RGB/Dep}_{IBP}+\mathcal{L}^{RGB/Dep}_{output}$ (6)
$\mathcal{L}^{RGB/Dep}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{IBP}^{Dep/RGB}}^{i})\}$ (7)
$\mathcal{L}^{RGB/Dep}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i})\}$ (8)
The overall loss $\mathcal{L_{CRD}}$ is expressed as
$\begin{split}\mathcal{L}_{CRD}&=(\mathcal{L}^{RGB}_{D}+\mathcal{L}^{RGB}_{CF}+\mathcal{L}^{RGB}_{CA})\\ &+(\mathcal{L}^{Dep}_{D}+\mathcal{L}^{Dep}_{CF}+\mathcal{L}^{Dep}_{CA})\end{split}$ (9)
Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.
4.4 Anomaly Scoring
Following previous KD-based methods, we generate the anomaly map for each branch using cosine distance as
$\mathcal{M}^{RGB/Dep}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i}))$ (10)
For the fusion of anomaly maps of two branches, we refer to the method in [8]:
  • First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{RGB/Dep}$ and $\sigma^{RGB/Dep}$, respectively.
  • Then, during inference, we normalize the anomaly maps for both branches as
    $\mathcal{M}^{RGB/Dep}_{norm}=\frac{\mathcal{M}^{RGB/Dep}-\mu^{RGB/Dep}}{\sigma^{RGB/Dep}}$ (11)
  • Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$.
And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
DepthGAN [7] 42.1 42.2 77.8 69.6 49.4 25.2 28.5 36.2 40.2 63.1 47.4
DepthAE [7] 43.2 15.8 80.8 49.1 84.1 40.6 26.2 21.6 71.6 47.8 48.1
DepthVM [7] 38.8 32.1 19.4 57.0 40.8 28.2 24.4 34.9 26.8 33.1 33.5
BTF [18] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
M3DM [28] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
Shape-guided [11] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
AST [24] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
MMRD [15] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
CFM [12] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
CRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
Table 2: Pixel-level anomaly localization results PRO (%) on MVTec 3D-AD with the best in bold.
5 Experiments
5.1 Experimental Settings
Dataset
We conduct experiments on MVTec 3D-AD [7] dataset which is a 3D industrial anomaly detection dataset that includes 10 categories of industrial objects. The training set contains a total of 2,656 normal samples, each consisting of an RGB image and corresponding 3D scan information. The test set includes 1,197 samples, comprising both 948 anomalous samples and 249 normal samples, and provides pixel-level annotations for anomaly detection and localization evaluation. Besides, a validation set is provided by MVTec 3D-AD, containing 294 normal samples.
Evaluation Metrics
The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUROC) and average precision (I-AP). For anomaly localization, pixel-level AUROC (P-AUROC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
Implementation Details
During both training and inference, only the depth information from the 3D data is used. All RGB images and depth maps are scaled to $256\times 256$. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet 50. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
Figure 4: Qualitative results on MVTec-3D AD. From top to bottom: the input images, the corresponding depth maps, the ground truth masks, the output anomaly maps of our proposed CRD.
5.2 Results on MVTec-3D AD
Tab. 1 and Tab. 2 show the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method, CRD, with several state-of-the-art 3D anomaly detection methods. CRD consistently outperforms all other methods in both I-AUROC and PRO metrics, achieving an average I-AUROC of 96.2% and an average PRO of 97.7%. Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach.
A subset of qualitative results on MVTec-3D AD is provided in Fig. 4, showing that our proposed CRD is able to accurately localize the anomaly regions. For further details on P-AUROC results and more qualitative results please refer to the Supplementary Materials.
5.3 Ablation Studies
Study on Multi-branch Distillation
Branch I-AUROC P-AUROC P-PRO I-AP P-AP
RGB 86.8 98.9 95.4 95.8 37.7
RGB+Crossmodal 88.9 99.0 95.8 96.5 39.4
Depth 87.1 96.2 88.8 95.2 33.7
Depth+Crossmodal 87.3 97.1 90.1 95.7 34.1
RGB+Depth 93.9 99.3 97.4 98 45.2
RGB+Depth+Crossmodal 96.2 99.4 97.7 98.8 45.5
Table 3: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
We conduct an ablation study in Tab. 3, which demonstrates that the multi-branch setup, combining both RGB and Depth modalities, significantly outperforms individual modality branches. The I-AUROC metric of 93.9% for using both branches is higher than 86.8% for RGB branch only and 87.1% for depth branch only, highlighting the complementary benefits of modality fusion.
Study on Crossmodal Assistants
CF CA I-AUROC P-AUROC P-PRO I-AP P-AP
- - 93.9 99.3 97.4 98 45.2
- 94.2 99.3 97.5 98.2 45.8
- 95.6 99.4 97.7 98.6 45.5
96.2 99.4 97.7 98.8 45.5
Table 4: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
We perform an ablation study to evaluate the impact of Crossmodal Filter and Crossmodal Amplifier on anomaly detection performance, as shown in Tab. 4. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves I-AUROC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUROC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUROC of 96.2%.
Study on Bottlneck Projection in Crossmodal Filter
Downsampling Size I-AUROC P-AUROC P-PRO I-AP P-AP
Identity 93.9 99.3 97.4 98.0 45.8
$16\times 16$ 93.6 99.3 97.4 97.9 45.2
$8\times 8$ 94.2 99.3 97.5 98.2 45.8
$4\times 4$ 93.8 99.3 97.4 98.0 45.3
Table 5: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab. 5. When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while too small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
Study on Inverted Bottlneck Projection in Crossmodal Amplifier
Channel Expansion I-AUROC P-AUROC P-PRO I-AP P-AP
Identity 95.7 99.3 97.7 98.7 45.1
2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5
4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0
Table 6: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
As recorded in Tab. 6, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier. The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency. This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
Study on Anomaly Scoring
Aggregation I-AUROC P-AUROC P-PRO I-AP P-AP
$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 94.3 99.3 97.2 98.2 43.5
Table 7: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
Tab. 7 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps. Normalization ensures the anomaly maps from different modalities are on the same scale, making fusion more effective. As a result, normalizing before aggregation improves performance across all metrics.
6 Conclusion
In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal industrial AD task by fully utilizing the information from all modalities. CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch. Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities. Experimental results on MVTec 3D-AD show that CRD achieves SOTA performance in both anomaly detection and localization.
Supplementary Material
Appendix A Motivation
Figure 5: Feature diagrams in the anomaly detection process based on Knowledge Distillation. The upper part illustrates previous methods (using RD framework as an example), while the lower part represents the feature transformation in our proposed method CRD. Green and orange indicate two modalities, while blue represents the fused results. In the resulting anomaly maps, darker gray regions indicate higher anomaly scores.
In Sec. 4.1, In Sec 4.1, we introduce the detection approach of previous unsupervised multimodal anomaly detection methods based on Knowledge Distillation. These methods detect and localize anomalies by analyzing the differences between the fused multimodal teacher features and student features. We also identify a key issue in these methods: the anomaly smoothing problem. When features from a normal modality dominate the fusion, abnormal information may be smoothed out, resulting in fused teacher features that fail to effectively reflect the anomalies.
This section provides an illustrative diagram to further explain this problem. In Fig. 5, the upper part outlines the feature transformation process of previous KD-based methods, while the lower part illustrates the feature transformation process in our proposed method. Unlike prior methods, which may miss anomalies when one modality is normal and the other is abnormal, our method ensures sensitivity to anomalies in each modality through dual-branch processing. Additionally, the proposed crossmodal assistants facilitate interaction between modalities, enabling more precise anomaly detection and localization.
Appendix B P-AUROC on on MVTec 3D-AD
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
BTF [18] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
M3DM [28] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
AST [24] - - - - - - - - - - 97.6
MMRD [15] - - - - - - - - - - 99.2
CFM [12] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
CRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
Table 8: Pixel-level anomaly localization results P-AUROC (%) on MVTec 3D-AD with the best in bold.
This section compares our method with other mainstream multimodal AD methods using the P-AUROC metric, including memory bank-based methods BTF [18] and M3DM [28], KD-based methods AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. As shown in Tab. 8, our method achieves SOTA performance, reaching 99.4% on the P-AUROC metric. Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods. In most other categories, it consistently maintains competitive or near-optimal performance.
Appendix C Additional Results of Ablation Studies
In the main text, we present five aspects of ablation studies:
  1. 1.
    Study on Multi-branch Distillation demonstrates that combining the distillation results from multiple branches of different modalities achieves better performance, highlighting the importance of multiple branches. (Tab. 9)
  2. 2.
    Study on Crossmodal Assistants shows that incorporating crossmodal assistants into Multi-branch Distillation based on Reverse Distillation significantly improves anomaly detection, proving their effectiveness. (Tab. 10)
  3. 3.
    Study on Bottleneck Projection in Crossmodal Filter explores the impact of downsampling sizes in the Bottleneck Projection module of Crossmodal Filter, concluding that a size of $8\times 8$ yields optimal results. (Tab. 11)
  4. 4.
    Study on Inverted Bottleneck Projection in Crossmodal Amplifier investigates the effect of channel expansion in the Inverted Bottleneck Projection module of Crossmodal Amplifier, finding that a 2×2$\times$2 × expansion achieves the best performance. (Tab. 12)
  5. 5.
    Study on Anomaly Scoring examines different methods for fusing anomaly maps from multiple branches, confirming that our selected approach delivers the best results. (Tab. 13)
In the main text, we provide average results across all categories for comparison. Here, we present detailed results for each category, as shown in Tab. 9, Tab. 10, Tab. 11, Tab. 12, and Tab. 13.
Metrics Branch Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUROC RGB 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
RGB+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
Depth 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
Depth+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
RGB+Depth 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
RGB+Depth+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUROC RGB 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
RGB+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
Depth 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
Depth+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
RGB+Depth 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
RGB+Depth+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO RGB 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
RGB+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
Depth 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
Depth+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
RGB+Depth 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
RGB+Depth+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP RGB 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
RGB+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
Depth 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
Depth+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
RGB+Depth 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
RGB+Depth+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP RGB 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
RGB+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
Depth 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
Depth+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
RGB+Depth 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
RGB+Depth+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
Table 9: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
Metrics CF CA Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUROC - - 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
- 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
- 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUROC - - 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
- 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
- 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO - - 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
- 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
- 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP - - 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
- 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
- 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP - - 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
- 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
- 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
Table 10: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
Metrics Downsampling Size Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUROC Identity 99.6 88.8 97.9 95.6 99.7 84.6 98.8 96.8 99.8 77.1 93.9
$16\times 16$ 99.7 88.1 97.6 96.9 100 81.6 99.2 96.7 99.8 76.7 93.6
$8\times 8$ 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
$4\times 4$ 99.7 89.5 98.1 95.6 99.9 81.6 98.8 96.8 99.8 77.5 93.8
P-AUROC Identity 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.1 99.3
$16\times 16$ 99.4 99.3 99.7 99.3 99.7 97.3 99.8 99.9 99.6 99.1 99.3
$8\times 8$ 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
$4\times 4$ 99.4 99.3 99.7 99.3 99.6 97.0 99.8 99.9 99.6 99.1 99.3
PRO Identity 98.2 98.0 99.0 98.1 98.4 90.2 99.3 99.5 98.0 95.6 97.4
$16\times 16$ 98.1 98.0 99.0 98.1 98.4 90.6 99.2 99.5 98.0 95.5 97.4
$8\times 8$ 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
$4\times 4$ 98.1 98.1 99.0 98.0 98.4 89.8 99.2 99.5 98.0 95.7 97.4
I-AP Identity 99.9 96.9 99.5 98.8 99.9 95.5 99.7 99.0 99.9 91.3 98.0
$16\times 16$ 99.9 96.7 99.5 99.2 100 93.9 99.8 98.9 99.9 90.9 97.9
$8\times 8$ 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
$4\times 4$ 99.9 97.3 99.6 98.8 100 93.8 99.7 98.9 99.9 91.8 98.0
P-AP Identity 45.7 37.5 41.2 69.3 40.6 27.8 62.2 57.5 52.8 23.2 45.8
$16\times 16$ 42.7 37.1 41.4 70.2 41.3 27.3 59.6 56.8 52.9 22.3 45.2
$8\times 8$ 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
$4\times 4$ 44.4 38.0 41.6 70.1 40.6 26.4 59.3 56.1 54.0 22.8 45.3
Table 11: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
Metrics Channel Expansion Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUROC Identity 99.6 95.2 98.3 98.0 99.7 87.4 99.2 97.6 100 81.9 95.7
2×2$\times$2 × 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
4×4$\times$4 × 99.5 94.7 98.0 98.6 99.7 88.8 99.2 97.5 100 80.8 95.7
P-AUROC Identity 99.3 99.4 99.7 99.3 99.7 97.4 99.8 99.5 99.7 99.4 99.3
2×2$\times$2 × 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
4×4$\times$4 × 99.3 99.3 99.7 99.3 99.7 97.6 99.8 99.9 99.7 99.4 99.4
PRO Identity 97.6 98.3 99.1 98.2 98.6 90.9 99.3 99.5 98.2 97.0 97.7
2×2$\times$2 × 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
4×4$\times$4 × 97.7 98.2 99.1 98.2 98.6 91.6 99.3 99.5 98.2 96.8 97.7
I-AP Identity 99.9 98.9 99.6 99.5 99.9 96.5 99.8 99.3 100 93.2 98.7
2×2$\times$2 × 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
4×4$\times$4 × 99.9 98.7 99.5 99.6 99.9 96.9 99.8 99.3 100 92.8 98.6
P-AP Identity 27.9 39.9 42.1 67.6 49.8 28.4 62.0 57.0 55.4 21.3 45.1
2×2$\times$2 × 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
4×4$\times$4 × 29.0 38.4 42.4 68.1 48.8 29.1 60.9 57.3 55.7 20.4 45.0
Table 12: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
Metrics Aggregation Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.6 92.9 98.3 98.3 99.3 90.3 98.9 96.7 100 81.6 95.6
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
P-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.4 96.8 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.1
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
PRO $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 97.8 96.8 99.1 98.2 98.7 91.1 99.3 99.5 98.2 97.2 97.6
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
I-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.9 98.2 99.6 99.6 99.8 97.2 99.7 99.0 100 92.7 98.6
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
P-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 31.6 27.0 43.9 70.4 48.8 29.6 60.3 55.9 55.3 21.2 44.4
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
Table 13: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
Appendix D Additional Qualitative Results
Figure 6: Qualitative results on MVTec-3D AD. From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.
This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig. 6. The visualizations reveal two key findings: (1) In the depth branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives. But this issue is less pronounced in the RGB branch. (2) Some anomalies are modality-specific. For instance, discoloration is only detectable in RGB images, while the absence of chocolate on cookies is only evident in depth maps. Consequently, anomaly maps from individual branches may fail to fully capture all anomalies within a sample. However, the fused results effectively achieve comprehensive anomaly localization.
Figure 7: Qualitative results on MVTec-3D AD. From top to bottom: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
We also provide visualizations of anomaly maps before and after incorporating the crossmodal assistants, as in Fig. 7. Without the crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing the crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
Appendix E Results on Eyecandies
E.1 Eyecandies Dataset
The Eyecandies dataset is a synthetic industrial dataset featuring 10 categories of candy samples in an industrial conveyor scenario. Each sample includes RGB images, depth map, and surface normal map. The training set contains 10k samples, with 1k samples per category. The validation and test sets comprise 1k validation samples and 4k test samples, respectively. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality for all samples.
E.2 Quantitative result
This section presents the quantitative results of anomaly detection and localization on Eyecandies in Tab. 14. The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [18] and M3DM [28], KD-based methods like AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. Our method significantly outperforms all non-KD-based multimodal anomaly detection methods. While it slightly lags behind MMRD [15] in anomaly detection metrics I-AUROC, it consistently achieves higher scores than all other methods including P-AUROC and PRO in anomaly localization.
Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
I-AUROC RGB-D [9] 52.9 86.1 73.9 75.2 59.4 49.8 67.9 65.1 83.8 75.0 68.9
RGB-cD-n [9] 59.6 84.3 81.9 84.6 83.3 55.0 75.0 84.6 94.0 84.8 78.7
M3DM [28] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
AST [24] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
MMRD [15] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
CFM [12] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
CRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
P-AUROC RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
M3DM [28] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [24] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [15] - - - - - - - - - - 98.3
CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
PRO M3DM [28] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
AST [24] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
MMRD [15] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
CFM [12] 94.2 90.2 83.1 96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
CRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
Table 14: Anomaly detection and localization results I-AUROC (%), P-AUROC (%), and PRO (%) on Eyecandies with the best in bold.
E.3 Qualitative result
Figure 8: Qualitative results on Eyecandies. From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.
Fig. 8 visualizes the anomaly maps generated by our proposed CRD on Eyecandies. Our method demonstrates strong sensitivity to both 2D and 3D modalities, enabling comprehensive multimodal anomaly localization.

### v2

@@
-secReferences
+Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation
 Abstract
-Knowledge distillation (KD) has been widely studied in unsupervised Industrial Image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored.
+Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored.
 Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges.
@@
 With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected.
-Experimental verifications on the MVTec 3D-AD dataset demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization.
-Code is available at CRD.
+Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization.
+Code will be available.
 1 Introduction
@@
 Given the limitations of RGB anomaly detection, data from 3D sensors, including depth map and point cloud, have been widely applied in AD in recent years, which results in growing research interest in unsupervised multimodal AD.
-Figure 1: Comparison of our proposed Crossmodal Reverse Distillation (CRD) with other multimodal AD methods based on feature learning.
-The upper part contains an overview of the two previous categories of methods, and the lower part shows the overall framework of CRD.
+Figure 1: Qualitative results on MVTec 3D-AD [7].
+From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [26] with more false negatives, Crossmodal Feature Mapping [12] (abbreviated as CFM) with more false positives, and our method CRD with superior segmentation.
+Figure 2: Comparison of our proposed CRD with other multimodal AD methods based on feature learning.
 Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [20] and M3DM [31].
-These methods construct large memory banks by processing and storing the multimodal features of normal samples from the training set.
-During inference, the extracted features are compared with the normal features in the memory bank to determine if the samples are anomalous.
+These methods construct large memory banks by processing and storing multimodal features of normal samples.
+During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous.
 However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
@@
 Currently, feature learning-based methods are divided into two categories as in Fig.
-1: (1) Methods like AST [24] and MMRD [15] (Fig.
+2: (1) Methods like AST [26] and MMRD [15] (Fig.
 2 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples.
-However, these methods may smooth out abnormal features during fusion, reducing teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous.
+However, these methods may smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous.
 (2) Crossmodal Feature Mapping [12] (Fig.
 2 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships.
-But inevitably, misalignment between multimodal features results in underfitting and cause false positives in the output anomaly maps.
+But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
 To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one.
@@
 Instead, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
-In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD), named Crossmodal Reverse Distillation (CRD).
+In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [13], named Crossmodal Reverse Distillation (CRD).
 CRD naturally combines the strengths of both feature learning paradigms.
-First, we design a Multi-branch Distillation (MBD) that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps.
+First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps.
 Next, we introduce two crossmodal assistants within each modality branch.
-One is Crossmodal Filter (CF), which uses information from other modalities to help the student decoder generate normal features.
-And the other is Crossmodal Amplifier (CA), which allows a modality branch to detect anomalies from other modalities.
-Finally, anomaly scores of the test samples are calculated by integrating the anomaly maps of all branches.
-Experimental results on the representative 3D AD dataset demonstrate the effectiveness of our method in multimodal anomaly detection and localization.
-Our contributions are as follows:
-  • We propose Crossmodal Reverse Distillation based on Multi-branch Distillation for multimodal industrial anomaly detection, which effectively detects anomalies in all modalities.
-  • Two crossmodal assistants are introduced including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
-  • Experimental results show that our proposed CRD achieves state-of-the-art performance on the multimodal industrial AD dataset.
+One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features.
+And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities.
+Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches.
+Experimental results on representative multimodal AD datasets demonstrate the effectiveness of CRD in multimodal anomaly detection and localization.
 2 Related Work
-Unsupervised Anomaly Detection
+Unsupervised Anomaly Detection.
 Unsupervised AD has gained increasing attention in recent years.
-Initially, many unsupervised AD methods relied on generative models [5, 1, 34].
+Initially, many unsupervised AD methods relied on generative models [5, 1, 38, 32, 19, 18].
 These models are trained on normal samples to learn the ability to reconstruct normal data.
@@
 Other methods introduce memory banks [25, 2, 22], comparing test samples with stored normal features to detect anomalies.
-Recently, there has been significant progress in the research of synthetic anomaly images [19, 30, 21], which simulate real-world scenarios to assist in unsupervised AD tasks.
-Additionally, knowledge distillation methods [25, 27, 24, 3, 13, 26, 16, 14, 33, 17] based on teacher-student framework have also been applied to unsupervised AD.
+Recently, significant progress has been made in the research of synthetic anomaly images [21, 34, 23], which simulate real-world scenarios to assist in unsupervised AD tasks.
+Additionally, knowledge distillation methods [27, 30, 26, 3, 13, 28, 16, 14, 37, 17] based on the teacher-student framework have also been applied to unsupervised AD.
 These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies.
 Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
-Multimodal Anomaly Detection
+Multimodal Anomaly Detection.
 With the release of multiple 3D industrial anomaly detection datasets [7, 9, 24], unsupervised multimodal AD has gradually become a topic of research.
-Some methods [11], such as BTF [18] and M3DM [28], follow PatchCore [23] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method.
+Some methods [11, 29], such as BTF [20] and M3DM [31], follow PatchCore [25] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method.
 Other methods [10, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data.
 Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [26], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network.
-Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two lightweight neural networks to perform crossmodal feature mapping and locate anomalies based on the mapping results.
+Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
 3 Preliminaries
@@
 The optimization loss is generally expressed as
-$\mathrm{Sim}(f_{1},f_{2})=\frac{{f_{1}}^{\mathsf{T}}f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
+$\mathrm{Sim}(f_{1},f_{2})=\frac{f_{1}\cdot f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
 $\mathcal{L}=\sum_{i=1}^{l}\{1-\mathrm{Sim}(F_{T}^{i},F_{S}^{i})\}$ (2)
 where $l$ represents the number of selected feature layers, typically set to 3.
-During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map, and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample.
+During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map.
+, and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample.
 When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low.
@@
 Unlike traditional forward distillation, RD uses an encoder-decoder architecture.
-The teacher network is a pre-trained encoder with a WideResNet 50 [29] backbone, while the student network is a trainable decoder that generates features matching the size of the teacher’s features.
-RD also includes a trainable bottleneck module, OCBE, to compress features between the teacher and student.
+The teacher network is a pre-trained encoder WideResNet-50 [33], while the student network is a trainable decoder that generates features matching the size of the teacher’s features.
+RD also includes a trainable one-class bottleneck embedding (OCBE) module to compress the teacher’s features brefore passing them to the student.
 During training, the student network learns to reconstruct the teacher’s normal features using normal samples.
@@
 At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network.
-As a result, even for abnormal samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
+Hence, even for anomalous samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
 4 Crossmodal Reverse Distillation
-Figure 2: Pipeline of our proposed method which introduces Knowledge Distillation to the multimodal industrial AD task.
-Our proposed method mainly consists of two parts: (1) Cross Reverse Distillation designs multiple branches to realize Reverse Distillation for multiple modalities, (2) Anomaly Scoring aggregates anomaly maps from multiple branches to achieve multimodal anomaly detection.
+Figure 3: Pipeline of our proposed method.
+The feature evolution visualization highlights that anomalies in each modality are effectively captured.
+In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
 Figure 4: Information flow of the branch of Modality A, showing how Crossmodal Filter and Crossmodal Amplifier incorporate information from Modality B.
 Crossmodal Filter aids the student decoder in generating normal features by utilizing features from Modality B, while Crossmodal Amplifier strengthens the perception of anomalies in Modality A by amplifying anomaly signals from Modality B.
-In the multimodal anomaly detection task, each input consists of multiple modalities of data.
-In this section, we primarily focus on RGB and Depth modalities as examples for studying multimodal anomaly detection.
-Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{RGB},\mathcal{I}^{Dep}\}$, where $\mathcal{I}^{RGB}$ and $\mathcal{I}^{Dep}$ are the RGB image and depth map respectively.
+In the multimodal AD task, each input consists of multiple modalities of data.
+In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection.
+Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively.
 The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$.
@@
 4.1 Multi-branch Distillation
-In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories:
-  • Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem.
-  • Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network.
+In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories: (1) Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem [26].
+(2) Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network [15].
 The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern.
@@
 To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig.
-2, which differs from previous KD-based multimodal AD methods with a multi-branch design.
-(In this section, we focus on the example case of dual-branch including RGB and depth branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality.
-During training, the pre-trained teacher network independently distills knowledge for each modality, eliminating the need for fusion at the input or feature exaction stages.
+3, which differs from previous KD-based multimodal AD methods with a multi-branch design.
+(In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality.
+During training, the pre-trained teacher network independently distills knowledge for each modality.
 At inference, an anomaly map is generated for each modality, and these maps are fused after normalization.
@@
 For the design of each branch, we draw inspiration from RD.
-Both the RGB and depth branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec.
+Both 2D and 3D branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec.
 4.2).
-The teacher encoders in both branches share parameters and are based on WideResNet 50 pre-trained on ImageNet.
-The student decoders follow the design of RD, using a structure symmetric to that of the teacher networks.
-Crossmodal Filter replaces the OCBE module of RD, while Crossmodal Amplifier is placed after the student decoder to optimize the output features, enabling crossmodal mapping and information fusion between modalities.
+The teacher encoders in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet.
+The student decoders follow RD, using a structure symmetric to that of the teacher networks.
+Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to optimize the output features, enabling crossmodal mapping and information fusion between modalities.
 The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
-Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-DepthGAN [7] 53.8 37.2 58.0 60.3 43.0 53.4 64.2 60.1 44.3 57.7 53.2
-DepthAE [7] 64.8 50.2 65.0 48.8 80.5 52.2 71.2 52.9 54.0 55.2 59.5
-DepthVM [7] 51.3 55.1 47.7 58.1 61.7 71.6 45.0 42.1 59.8 62.3 55.5
-BTF [18] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
-M3DM [28] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
-Shape-guided [11] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
-AST [24] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
-MMRD [15] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
-CFM [12] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
-CRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-Table 1: Image-level anomaly detection results I-AUROC (%) on MVTec 3D-AD with the best in bold. We abbreviate Crossmodal Feature Mapping as CFM, the same below.
 4.2 Crossmodal Assistants
@@
 (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality.
-With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final detection.
-Crossmodal Filter
+With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
+Crossmodal Filter.
 The training objective of RD is to ensure that the student decoder generates anomaly-free features.
@@
 During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
-Crossmodal Amplifier
+Crossmodal Amplifier.
 When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec.
@@
 As illustrated in Fig.
-3, we design Crossmodal Amplifier.
+4, we design Crossmodal Amplifier.
 First, Inverted Bottleneck Projection (IBP) is designed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$.
@@
 In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes.
-We believe that this process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information in the modality during feature mapping.
+This process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information during feature mapping.
+Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Avg.
+DepthGAN [7] 53.8/42.1 37.2/42.2 58.0/77.8 60.3/69.6 43.0/49.4 53.4/25.2 64.2/28.5 60.1/36.2 44.3/40.2 57.7/63.1 53.2/47.4
+DepthAE [7] 64.8/43.2 50.2/15.8 65.0/80.8 48.8/49.1 80.5/84.1 52.2/40.6 71.2/26.2 52.9/21.6 54.0/71.6 55.2/47.8 59.5/48.1
+DepthVM [7] 51.3/38.8 55.1/32.1 47.7/19.4 58.1/57.0 61.7/40.8 71.6/28.2 45.0/24.4 42.1/34.9 59.8/26.8 62.3/33.1 55.5/33.5
+BTF [20] 91.8/97.6 74.8/96.9 96.7/97.9 88.3/97.3 93.2/93.3 58.2/88.8 89.6/97.5 91.2/98.1 92.1/95.0 88.6/97.1 86.5/95.9
+M3DM [31] 99.4/97.0 90.9/97.1 97.2/97.9 97.6/95.0 96.0/94.1 94.2/93.2 97.3/97.7 89.9/97.1 97.2/97.1 85.0/97.5 94.5/96.4
+Shape-guided [11] 98.6/98.1 89.4/97.3 98.3/98.2 99.1/97.1 97.6/96.2 85.7/97.8 99.0/98.1 96.5/98.3 96.0/97.4 86.9/97.5 94.7/97.6
+AST [26] 98.3/97.0 87.3/94.7 97.6/98.1 97.1/93.9 93.2/91.3 88.5/90.6 97.4/97.9 98.1/98.2 100/88.9 79.7/94.0 93.7/94.4
+MMRD [15] 99.9/98.6 94.3/99.0 96.4/99.1 94.3/95.1 99.2/99.0 91.2/90.1 94.9/99.0 90.1/99.0 99.4/98.7 90.1/98.2 95.0/97.6
+CFM [12] 98.8/98.0 87.5/96.6 98.4/98.2 99.2/94.7 99.7/95.9 92.4/96.7 96.4/98.2 94.9/98.3 97.9/97.6 95.0/98.2 96.0/97.2
+CRD (Ours) 99.6/97.7 96.1/98.1 98.2/99.1 98.5/98.2 99.7/98.7 90.1/91.0 99.2/99.4 97.6/99.5 100/98.2 83.3/97.4 96.2/97.7
+Table 1: Anomaly detection and localization results I-AUROC/PRO (%) on MVTec 3D-AD with the best in bold and the second-best underlined.
 4.3 Training Objectives
@@
 Firstly, similar to the approach in RD, we optimize the features output by the student decoder using cosine similarity as
-$\mathcal{L}^{RGB/Dep}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{D}^{RGB/Dep}}^{i})\}$ (4)
-where $\mathcal{L}^{RGB}_{D}$ and $\mathcal{L}^{Dep}_{D}$ are the losses corresponding to RGB and depth branches respectively.
-Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{Dep/RGB}$ within these filters to the encoder features.
-The training losses of RGB and Depth branches are calculated as:
-$\mathcal{L}^{RGB/Dep}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{BP}^{Dep/RGB}}^{i})\}$ (5)
+$\mathcal{L}^{2D/3D}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{D}^{2D/3D}}^{i})\}$ (4)
+where $\mathcal{L}^{2D}_{D}$ and $\mathcal{L}^{3D}_{D}$ are the losses corresponding to 2D and 3D branches respectively.
+Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ within these filters to the encoder features.
+The training losses of 2D and 3D branches are calculated as:
+$\mathcal{L}^{2D/3D}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{BP}^{3D/2D}}^{i})\}$ (5)
 Finally, Crossmodal Amplifier needs to be optimized,, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. The total loss function is
-$\mathcal{L}^{RGB/Dep}_{CA}=\mathcal{L}^{RGB/Dep}_{IBP}+\mathcal{L}^{RGB/Dep}_{output}$ (6)
-$\mathcal{L}^{RGB/Dep}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{IBP}^{Dep/RGB}}^{i})\}$ (7)
-$\mathcal{L}^{RGB/Dep}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i})\}$ (8)
+$\mathcal{L}^{2D/3D}_{CA}=\mathcal{L}^{2D/3D}_{IBP}+\mathcal{L}^{2D/3D}_{output}$ (6)
+$\mathcal{L}^{2D/3D}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{IBP}^{3D/2D}}^{i})\}$ (7)
+$\mathcal{L}^{2D/3D}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i})\}$ (8)
 The overall loss $\mathcal{L}_{CRD}$ is expressed as
-$\begin{split}\mathcal{L}_{CRD}&=(\mathcal{L}^{RGB}_{D}+\mathcal{L}^{RGB}_{CF}+\mathcal{L}^{RGB}_{CA})\\ &+(\mathcal{L}^{Dep}_{D}+\mathcal{L}^{Dep}_{CF}+\mathcal{L}^{Dep}_{CA})\end{split}$ (9)
+$\mathcal{L}_{CRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
 Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.
@@
 Following previous KD-based methods, we generate the anomaly map for each branch using cosine distance as
-$\mathcal{M}^{RGB/Dep}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{RGB/Dep}}^{i},{F_{CA}^{RGB/Dep}}^{i}))$ (10)
-For the fusion of anomaly maps of two branches, we refer to the method in [8]:
-  • First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{RGB/Dep}$ and $\sigma^{RGB/Dep}$, respectively.
+$\mathcal{M}^{2D/3D}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i}))$ (10)
+For the fusion of anomaly maps of two branches, we refer to the method in [8]: First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{2D/3D}$ and $\sigma^{2D/3D}$, respectively.
 Then, during inference, we normalize the anomaly maps for both branches as
-    $\mathcal{M}^{RGB/Dep}_{norm}=\frac{\mathcal{M}^{RGB/Dep}-\mu^{RGB/Dep}}{\sigma^{RGB/Dep}}$ (11)
-  • Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$.
+$\mathcal{M}^{2D/3D}_{norm}=\frac{\mathcal{M}^{2D/3D}-\mu^{2D/3D}}{\sigma^{2D/3D}}$ (11)
+Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{2D}_{norm}+\mathcal{M}^{3D}_{norm}$.
 And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
-Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-DepthGAN [7] 42.1 42.2 77.8 69.6 49.4 25.2 28.5 36.2 40.2 63.1 47.4
-DepthAE [7] 43.2 15.8 80.8 49.1 84.1 40.6 26.2 21.6 71.6 47.8 48.1
-DepthVM [7] 38.8 32.1 19.4 57.0 40.8 28.2 24.4 34.9 26.8 33.1 33.5
-BTF [18] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
-M3DM [28] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
-Shape-guided [11] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
-AST [24] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
-MMRD [15] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
-CFM [12] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
-CRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
-Table 2: Pixel-level anomaly localization results PRO (%) on MVTec 3D-AD with the best in bold.
+Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
+M3DM [31] 62.4/90.6 95.8/92.3 95.8/80.3 100/98.3 88.6/85.5 75.8/68.8 94.9/88.0 83.6/90.6 100/96.6 100/95.5 89.7/88.2
+AST [26] 57.4/51.4 74.7/83.5 74.7/71.4 88.9/90.5 59.6/58.7 61.7/59.0 81.6/73.6 84.1/76.9 98.7/91.8 98.7/87.8 78.0/74.4
+MMRD [15] 85.4/97.5 100/97.0 94.6/94.2 99.8/98.5 90.8/91.7 74.7/68.0 96.6/97.0 98.4/94.1 100/99.0 100/99.2 94.0/93.6
+CFM [12] 68.0/94.2 93.1/90.2 95.2/83.1 88.0/96.5 86.5/87.5 78.2/76.2 91.7/79.1 84.0/91.3 99.8/93.9 96.2/94.9 88.1/88.7
+CRD (Ours) 86.7/98.0 100/94.1 98.1/96.8 98.1/98.1 87.2/93.6 72.8/77.3 97.1/98.1 93.0/94.4 100/98.9 99.5/98.1 93.3/94.7
+Table 2: Anomaly detection and localization results I-AUC/PRO (%) on Eyecandies with the best in bold and the second-best underlined.
 5 Experiments
 5.1 Experimental Settings
-Dataset
-We conduct experiments on MVTec 3D-AD [7] dataset which is a 3D industrial anomaly detection dataset that includes 10 categories of industrial objects.
-The training set contains a total of 2,656 normal samples, each consisting of an RGB image and corresponding 3D scan information.
-The test set includes 1,197 samples, comprising both 948 anomalous samples and 249 normal samples, and provides pixel-level annotations for anomaly detection and localization evaluation.
-Besides, a validation set is provided by MVTec 3D-AD, containing 294 normal samples.
-Evaluation Metrics
-The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUROC) and average precision (I-AP).
-For anomaly localization, pixel-level AUROC (P-AUROC) and AP (P-AP) are used.
+Dataset.
+We conduct experiments MVTec 3D-AD [7] and Eyecandies [9].
+MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects.
+The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples.
+The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation.
+For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality.
+Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps.
+It contains 10k training samples, 1k validation samples, and 4k test samples.
+In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
+Evaluation Metrics.
+The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP).
+For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used.
 We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
-Implementation Details
-During both training and inference, only the depth information from the 3D data is used.
-All RGB images and depth maps are scaled to $256\times 256$.
+Implementation Details.
+All RGB images, depth maps and normal maps are scaled to $256\times 256$.
 The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category.
-The network backbone is WideResNet 50.
+The network backbone is WideResNet-50 [33].
 The batch size is 16, and the training lasts for 200 epochs.
@@
 In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
-Figure 4: Qualitative results on MVTec-3D AD. From top to bottom: the input images, the corresponding depth maps, the ground truth masks, the output anomaly maps of our proposed CRD.
-5.2 Results on MVTec-3D AD
+Figure 5: Qualitative ablation results of crossmodal assistants on MVTec 3D-AD.
+5.2 Main Results
+Results on MVTec 3D-AD.
 Tab.
-1 and Tab.
-2 show the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset.
-We compare our method, CRD, with several state-of-the-art 3D anomaly detection methods.
-CRD consistently outperforms all other methods in both I-AUROC and PRO metrics, achieving an average I-AUROC of 96.2% and an average PRO of 97.7%.
+1 shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset.
+We compare our method CRD with several state-of-the-art 3D anomaly detection methods.
+CRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%.
 Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach.
-A subset of qualitative results on MVTec-3D AD is provided in Fig.
-4, showing that our proposed CRD is able to accurately localize the anomaly regions.
-For further details on P-AUROC results and more qualitative results please refer to the Supplementary Materials.
+A subset of qualitative results on MVTec 3D-AD is provided in Fig.
+1, showing that our proposed CRD accurately localizes the anomaly regions.
+Results on Eyecandies.
+We evaluate our method on the Eyecandies dataset, as shown in Tab.
+2.
+Our method CRD significantly outperforms all non-KD-based multimodal anomaly detection methods.
+While it slightly lags behind MMRD [15] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
+Method MVTec 3D-AD (%) Eyecandies (%) Avg. (%) Frame Rate (fps)
+I-AUC P-AUC PRO I-AUC P-AUC PRO
+AST [26] 93.7 97.6 94.4 78.0 90.2 74.4 88.1 13.9
+MMRD* [15] 95.0 99.2 97.6 94.0 98.3 93.6 96.3 -
+CFM [12] 96.0 99.3 97.2 88.1 97.4 88.7 94.5 12.0
+CRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
+Table 3: Comparison of feature learning-based methods on MVTec 3D-AD and Eyecandies datasets under the same hardware setup. * indicates that MMRD has no reproducible code.
+Comparison with Feature Learning-based Methods.
+We further compare CRD with feature learning-based methods including AST[26], MMRD[15], and CFM [12] on MVTec 3D-AD and Eyecandies.
+As in Table 3, CRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps.
+Despite the multi-branch design, CRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
 5.3 Ablation Studies
-Study on Multi-branch Distillation
-Branch I-AUROC P-AUROC P-PRO I-AP P-AP
-RGB 86.8 98.9 95.4 95.8 37.7
-RGB+Crossmodal 88.9 99.0 95.8 96.5 39.4
-Depth 87.1 96.2 88.8 95.2 33.7
-Depth+Crossmodal 87.3 97.1 90.1 95.7 34.1
-RGB+Depth 93.9 99.3 97.4 98 45.2
-RGB+Depth+Crossmodal 96.2 99.4 97.7 98.8 45.5
+Branch I-AUC P-AUC P-PRO I-AP P-AP Avg.
+2D 86.8 98.9 95.4 95.8 37.7 82.9
+2D+Crossmodal 88.9 99.0 95.8 96.5 39.4 83.9
+3D 87.1 96.2 88.8 95.2 33.7 80.2
+3D+Crossmodal 87.3 97.1 90.1 95.7 34.1 80.9
+2D+3D 93.9 99.3 97.4 98.0 45.2 86.8
+2D+3D+Crossmodal 96.2 99.4 97.7 98.8 45.5 87.5
 Table 4: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
+CF CA I-AUC P-AUC P-PRO I-AP P-AP Avg.
+- - 93.9 99.3 97.4 98.0 45.2 86.8
+- 94.2 99.3 97.5 98.2 45.8 87.0
+- 95.6 99.4 97.7 98.6 45.5 87.4
+96.2 99.4 97.7 98.8 45.5 87.5
+Table 5: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
+Study on Multi-branch Distillation.
 We conduct an ablation study in Tab.
-3, which demonstrates that the multi-branch setup, combining both RGB and Depth modalities, significantly outperforms individual modality branches.
-The I-AUROC metric of 93.9% for using both branches is higher than 86.8% for RGB branch only and 87.1% for depth branch only, highlighting the complementary benefits of modality fusion.
-Study on Crossmodal Assistants
-CF CA I-AUROC P-AUROC P-PRO I-AP P-AP
-- - 93.9 99.3 97.4 98 45.2
-- 94.2 99.3 97.5 98.2 45.8
-- 95.6 99.4 97.7 98.6 45.5
-96.2 99.4 97.7 98.8 45.5
-Table 4: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
-We perform an ablation study to evaluate the impact of Crossmodal Filter and Crossmodal Amplifier on anomaly detection performance, as shown in Tab.
-4.
+4, which demonstrates that the multi-branch setup, combining both 2D and 3D modalities, significantly outperforms individual modality branches.
+The I-AUC metric of 93.9% for using both branches is higher than 86.8% for 2D branch only and 87.1% for 3D branch only, highlighting the complementary benefits of modality fusion.
+Study on Crossmodal Assistants.
+We perform an ablation study to evaluate the impact of crossmodal assistants on anomaly detection performance, as shown in Tab.
+5.
 The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results.
-Specifically, adding CF to the baseline improves I-AUROC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUROC of 95.6%.
-The highest performance is achieved when both CF and CA are used together, resulting in an I-AUROC of 96.2%.
-Study on Bottlneck Projection in Crossmodal Filter
-Downsampling Size I-AUROC P-AUROC P-PRO I-AP P-AP
-Identity 93.9 99.3 97.4 98.0 45.8
-$16\times 16$ 93.6 99.3 97.4 97.9 45.2
-$8\times 8$ 94.2 99.3 97.5 98.2 45.8
-$4\times 4$ 93.8 99.3 97.4 98.0 45.3
+Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%.
+The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
+We also provide visualizations of anomaly maps before and after incorporating crossmodal assistants, as in Fig.
+5.
+Without crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch.
+However, after introducing crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
+Downsampling Size I-AUC P-AUC P-PRO I-AP P-AP Avg.
+Identity 93.9 99.3 97.4 98.0 45.8 86.9
+$16\times 16$ 93.6 99.3 97.4 97.9 45.2 86.7
+$8\times 8$ 94.2 99.3 97.5 98.2 45.8 87.0
+$4\times 4$ 93.8 99.3 97.4 98.0 45.3 86.8
 Table 6: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
+Study on Bottlneck Projection in Crossmodal Filter.
 We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab.
-5.
+6.
 When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics.
-It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while too small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively.
+It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively.
 Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
-Study on Inverted Bottlneck Projection in Crossmodal Amplifier
-Channel Expansion I-AUROC P-AUROC P-PRO I-AP P-AP
-Identity 95.7 99.3 97.7 98.7 45.1
-2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5
-4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0
+Channel Expansion I-AUC P-AUC P-PRO I-AP P-AP Avg.
+Identity 95.7 99.3 97.7 98.7 45.1 87.3
+2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5 87.4
+4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0 87.3
 Table 7: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
+Study on Inverted Bottlneck Projection in Crossmodal Amplifier.
 As recorded in Tab.
-6, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier.
+7, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier.
 The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency.
 This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
-Study on Anomaly Scoring
-Aggregation I-AUROC P-AUROC P-PRO I-AP P-AP
-$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 94.3 99.3 97.2 98.2 43.5
+Aggregation I-AUC P-AUC P-PRO I-AP P-AP Avg.
+$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5 87.5
+$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4 87.1
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 94.3 99.3 97.2 98.2 43.5 86.5
 Table 8: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
+Study on Anomaly Scoring.
 Tab.
 8 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps.
-Normalization ensures the anomaly maps from different modalities are on the same scale, making fusion more effective.
-As a result, normalizing before aggregation improves performance across all metrics.
+Normalization ensures that anomaly maps from different modalities are on the same scale, making fusion more effective.
+Figure 6: Visualization of the features output by CF and CA.
+Visualization of Mapped Features.
+We visualizes the output features of CF and CA using the visualization decoders [32] in Fig.
+6, which intuitively demonstrates the roles of CA and CF.
+Regardless of whether the initial modality is normal or anomalous, CF consistently produces normal outputs, enabling the student decoder of the other modality to generate anomaly-free results.
+When the initial modality is anomalous, CA introduces discrepancies with the other modality, amplifying anomalies for enhanced detection.
 6 Conclusion
-In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal industrial AD task by fully utilizing the information from all modalities.
+In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task.
 CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch.
 Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities.
-Experimental results on MVTec 3D-AD show that CRD achieves SOTA performance in both anomaly detection and localization.
+Experimental results show that CRD achieves SOTA performance in both anomaly detection and localization.
 Supplementary Material
-Appendix A Motivation
-Figure 5: Feature diagrams in the anomaly detection process based on Knowledge Distillation.
-The upper part illustrates previous methods (using RD framework as an example), while the lower part represents the feature transformation in our proposed method CRD.
-Green and orange indicate two modalities, while blue represents the fused results.
-In the resulting anomaly maps, darker gray regions indicate higher anomaly scores.
-In Sec.
-4.1, In Sec 4.1, we introduce the detection approach of previous unsupervised multimodal anomaly detection methods based on Knowledge Distillation.
-These methods detect and localize anomalies by analyzing the differences between the fused multimodal teacher features and student features.
-We also identify a key issue in these methods: the anomaly smoothing problem.
-When features from a normal modality dominate the fusion, abnormal information may be smoothed out, resulting in fused teacher features that fail to effectively reflect the anomalies.
-This section provides an illustrative diagram to further explain this problem.
-In Fig.
-5, the upper part outlines the feature transformation process of previous KD-based methods, while the lower part illustrates the feature transformation process in our proposed method.
-Unlike prior methods, which may miss anomalies when one modality is normal and the other is abnormal, our method ensures sensitivity to anomalies in each modality through dual-branch processing.
-Additionally, the proposed crossmodal assistants facilitate interaction between modalities, enabling more precise anomaly detection and localization.
-Appendix B P-AUROC on on MVTec 3D-AD
+Overview
+This supplementary material provides extended analyses and visualizations to complement the main paper:
+  • Detailed P-AUC results for anomaly localization on MVTec 3D-AD and Eyecandies datasets. (Appendix A)
+  • Training accuracy curves to demonstrate the model’s convergence over training. (Appendix B)
+  • Detailed results of ablation studies on MVTec 3D-AD dataset. (Appendix C)
+  • Additional qualitative visualizations on MVTec 3D-AD and Eyecandies datasets. (Appendix D)
+Appendix A P-AUC Metric
+A.1 P-AUC on MVTec 3D-AD
 Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
@@
 CRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-Table 8: Pixel-level anomaly localization results P-AUROC (%) on MVTec 3D-AD with the best in bold.
-This section compares our method with other mainstream multimodal AD methods using the P-AUROC metric, including memory bank-based methods BTF [18] and M3DM [28], KD-based methods AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
+Table 9: Pixel-level anomaly localization results P-AUC (%) on MVTec 3D-AD with the best in bold.
+This section compares our method with other mainstream multimodal AD methods on MVTec 3D-AD dataset using the P-AUC metric, including memory bank-based methods BTF [20] and M3DM [31], KD-based methods AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
 As shown in Tab.
-8, our method achieves SOTA performance, reaching 99.4% on the P-AUROC metric.
+9, our method achieves SOTA performance, reaching 99.4% on the P-AUC metric.
 Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods.
 In most other categories, it consistently maintains competitive or near-optimal performance.
+A.2 P-AUC on Eyecandies
+Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
+RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
+RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
+M3DM [31] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
+AST [26] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
+MMRD [15] - - - - - - - - - - 98.3
+CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
+CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
+Table 10: Pixel-level anomaly localization results P-AUC (%) on Eyecandies with the best in bold.
+This section presents the quantitative results of anomaly localization on Eyecandies dataset using the P-AUC metric in Tab.
+10.
+The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [20] and M3DM [31], KD-based methods like AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
+Our method significantly outperforms all non-KD-based multimodal anomaly detection methods, achieving 98.5%.
+Appendix B Training Curves: I-AUC, P-AUC, and PRO
+(a) I-AUC.
+(b) P-AUC.
+(c) PRO.
+Figure 7: AD performance over 500 training epochs.
+We provide training curves of I-AUC, P-AUC, and PRO for the carrot and peach categories from the MVTec 3D-AD dataset.
+As shown in Fig.
+7, all metrics stabilize after approximately 50 epochs, indicating that the model converges efficiently without significant performance fluctuations.
+This case demonstrates the robustness of our method, as it avoids over-generalization even with extended training.
+Following the setting of Reverse Distillation [13], we adopt 200 epochs for training to ensure sufficient convergence while maintaining computational efficiency.
 Appendix C Additional Results of Ablation Studies
@@
 (Tab.
-9)
+11)
   2. 2.
@@
 (Tab.
-10)
+12)
   3. 3.
@@
 (Tab.
-11)
+13)
   4. 4.
@@
 (Tab.
-12)
+14)
   5. 5.
@@
 Metrics Branch Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUROC RGB 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
-RGB+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
-Depth 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
-Depth+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
-RGB+Depth 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
-RGB+Depth+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-P-AUROC RGB 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
-RGB+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
-Depth 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
-Depth+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
-RGB+Depth 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
-RGB+Depth+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-PRO RGB 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
-RGB+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
-Depth 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
-Depth+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
-RGB+Depth 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
-RGB+Depth+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
-I-AP RGB 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
-RGB+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
-Depth 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
-Depth+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
-RGB+Depth 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
-RGB+Depth+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
-P-AP RGB 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
-RGB+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
-Depth 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
-Depth+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
-RGB+Depth 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
-RGB+Depth+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
+I-AUC 2D 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
+2D+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
+3D 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
+3D+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
+2D+3D 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
+2D+3D+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
+P-AUC 2D 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
+2D+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
+3D 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
+3D+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
+2D+3D 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
+2D+3D+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
+PRO 2D 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
+2D+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
+3D 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
+3D+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
+2D+3D 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
+2D+3D+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
+I-AP 2D 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
+2D+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
+3D 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
+3D+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
+2D+3D 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
+2D+3D+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
+P-AP 2D 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
+2D+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
+3D 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
+3D+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
+2D+3D 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
+2D+3D+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
 Table 11: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
 Metrics CF CA Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUROC - - 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
+I-AUC - - 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
 - 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
@@
 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-P-AUROC - - 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
+P-AUC - - 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
 - 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
@@
 Metrics Downsampling Size Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUROC Identity 99.6 88.8 97.9 95.6 99.7 84.6 98.8 96.8 99.8 77.1 93.9
+I-AUC Identity 99.6 88.8 97.9 95.6 99.7 84.6 98.8 96.8 99.8 77.1 93.9
 $16\times 16$ 99.7 88.1 97.6 96.9 100 81.6 99.2 96.7 99.8 76.7 93.6
@@
 $4\times 4$ 99.7 89.5 98.1 95.6 99.9 81.6 98.8 96.8 99.8 77.5 93.8
-P-AUROC Identity 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.1 99.3
+P-AUC Identity 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.1 99.3
 $16\times 16$ 99.4 99.3 99.7 99.3 99.7 97.3 99.8 99.9 99.6 99.1 99.3
@@
 Metrics Channel Expansion Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUROC Identity 99.6 95.2 98.3 98.0 99.7 87.4 99.2 97.6 100 81.9 95.7
+I-AUC Identity 99.6 95.2 98.3 98.0 99.7 87.4 99.2 97.6 100 81.9 95.7
 2×2$\times$2 × 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
 4×4$\times$4 × 99.5 94.7 98.0 98.6 99.7 88.8 99.2 97.5 100 80.8 95.7
-P-AUROC Identity 99.3 99.4 99.7 99.3 99.7 97.4 99.8 99.5 99.7 99.4 99.3
+P-AUC Identity 99.3 99.4 99.7 99.3 99.7 97.4 99.8 99.5 99.7 99.4 99.3
 2×2$\times$2 × 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
@@
 Metrics Aggregation Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
+I-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
 $\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.6 92.9 98.3 98.3 99.3 90.3 98.9 96.7 100 81.6 95.6
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
-P-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
+P-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
 $\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.4 96.8 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.1
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
 PRO $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
 $\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 97.8 96.8 99.1 98.2 98.7 91.1 99.3 99.5 98.2 97.2 97.6
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
 I-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
 $\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.9 98.2 99.6 99.6 99.8 97.2 99.7 99.0 100 92.7 98.6
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
 P-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
 $\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 31.6 27.0 43.9 70.4 48.8 29.6 60.3 55.9 55.3 21.2 44.4
-$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
+$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
 Table 15: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
 Appendix D Additional Qualitative Results
-Figure 6: Qualitative results on MVTec-3D AD.
-From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.
+Figure 8: Qualitative results on MVTec 3D-AD.
+From top to bottom: 2D images, 3D depth maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
+D.1 Qualitative Results on MVTec 3D-AD
 This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig.
-6.
-The visualizations reveal two key findings: (1) In the depth branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives.
-But this issue is less pronounced in the RGB branch.
+8.
+The visualizations reveal two key findings: (1) In the 3D branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives.
+But this issue is less pronounced in the 2D branch.
 (2) Some anomalies are modality-specific.
@@
 However, the fused results effectively achieve comprehensive anomaly localization.
-Figure 7: Qualitative results on MVTec-3D AD. From top to bottom: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
-We also provide visualizations of anomaly maps before and after incorporating the crossmodal assistants, as in Fig.
-7.
-Without the crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch.
-However, after introducing the crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
-Appendix E Results on Eyecandies
-E.1 Eyecandies Dataset
-The Eyecandies dataset is a synthetic industrial dataset featuring 10 categories of candy samples in an industrial conveyor scenario.
-Each sample includes RGB images, depth map, and surface normal map.
-The training set contains 10k samples, with 1k samples per category.
-The validation and test sets comprise 1k validation samples and 4k test samples, respectively.
-In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality for all samples.
-E.2 Quantitative result
-This section presents the quantitative results of anomaly detection and localization on Eyecandies in Tab.
-14.
-The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [18] and M3DM [28], KD-based methods like AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
-Our method significantly outperforms all non-KD-based multimodal anomaly detection methods.
-While it slightly lags behind MMRD [15] in anomaly detection metrics I-AUROC, it consistently achieves higher scores than all other methods including P-AUROC and PRO in anomaly localization.
-Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
-I-AUROC RGB-D [9] 52.9 86.1 73.9 75.2 59.4 49.8 67.9 65.1 83.8 75.0 68.9
-RGB-cD-n [9] 59.6 84.3 81.9 84.6 83.3 55.0 75.0 84.6 94.0 84.8 78.7
-M3DM [28] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
-AST [24] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
-MMRD [15] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
-CFM [12] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
-CRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
-P-AUROC RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
-RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
-M3DM [28] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
-AST [24] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
-MMRD [15] - - - - - - - - - - 98.3
-CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
-CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
-PRO M3DM [28] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
-AST [24] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
-MMRD [15] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
-CFM [12] 94.2 90.2 83.1 96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
-CRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
-Table 14: Anomaly detection and localization results I-AUROC (%), P-AUROC (%), and PRO (%) on Eyecandies with the best in bold.
-E.3 Qualitative result
+Figure 9: Qualitative results on MVTec 3D-AD. From left to right: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
+In the main text, we visualize anomaly maps with and without crossmodal assistants, demonstrating their effectiveness in improving anomaly segmentation accuracy by reducing false positives and false negatives.
+In this section, we provide additional examples to further validate the robustness of our method.
+As shown in Fig.
+9, the inclusion of CF and CA consistently enhances the precision of anomaly localization across various categories, including bagel, cable cland, cookie, dowel, and potato.
+These results highlight the critical role of crossmodal interactions in refining anomaly detection.
+D.2 Qualitative result on Eyecandies
 Figure 10: Qualitative results on Eyecandies.
-From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.
+From top to bottom: the input 2D images, the corresponding 3D normal maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
 Fig.

### v3

@@
-Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation
+Tuned Reverse Distillation: Enhancing Multimodal Industrial Anomaly Detection with Crossmodal Tuners
 Abstract
@@
 Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges.
-Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures.
-Besides, these methods do not fully leverage the rich intra- and inter-modality information.
-In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD.
+Anomalies that only exist in one modality may not be effectively captured in the fused teacher features, leading to detection failures.
+Besides, these methods do not fully leverage the rich intra- and inter-modality information that are critical for effective anomaly detection.
+In this paper, we propose Tuned Reverse Distillation (TRD) based on Multi-branch design to realize Multimodal Industrial AD.
 By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality.
-Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier.
+Furthermore, we enhance the interaction between modalities during the distillation process by designing two Crossmodal Tuners including Crossmodal Filter and Amplifier.
 With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected.
 Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization.
-Code will be available.
-1 Introduction
+Code is available at https://github.com/hito2448/TRD.
+Index Terms:
+Multimodal anomaly detection, knowledge distillation, unsupervised anomaly detection, anomaly localization.
+I Introduction
 Industrial Anomaly Detection (AD) is an important task aimed at ensuring product quality by identifying defects in products.
@@
 Given the limitations of RGB anomaly detection, data from 3D sensors, including depth map and point cloud, have been widely applied in AD in recent years, which results in growing research interest in unsupervised multimodal AD.
-Figure 1: Qualitative results on MVTec 3D-AD [7].
-From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [26] with more false negatives, Crossmodal Feature Mapping [12] (abbreviated as CFM) with more false positives, and our method CRD with superior segmentation.
-Figure 2: Comparison of our proposed CRD with other multimodal AD methods based on feature learning.
-Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [20] and M3DM [31].
+Figure 1: Qualitative comparisons on MVTec 3D-AD [4].
+From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [5] with more false negatives, Crossmodal Feature Mapping [6] (abbreviated as CFM) with more false positives, and our method TRD with superior segmentation.
+Early unsupervised multimodal AD methods often rely on memory banks, such as BTF [7] and M3DM [8].
 These methods construct large memory banks by processing and storing multimodal features of normal samples.
@@
 Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [5, 9, 6].
-The feature learning methods typically leverage pretrained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks.
+The feature learning-based methods typically leverage pre-trained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks.
 Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks.
@@
 During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
-Currently, feature learning-based methods are divided into two categories as in Fig.
-2: (1) Methods like AST [26] and MMRD [15] (Fig.
-2 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples.
-However, these methods may smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous.
-(2) Crossmodal Feature Mapping [12] (Fig.
-2 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships.
+Figure 2: Comparison of our proposed TRD with previous KD-based methods.
+Previous KD-based methods tend to suffer from anomaly smoothing when one modality is normal and the other contains anomalies.
+For example, in this figure, the anomaly on the bagel is barely visible in the 2D modality and resembles normal samples, while it is clearly distinguishable in the 3D modality.
+However, one of previous KD-based methods AST [5], which fuses the two modalities before distillation, fails to detect the anomaly and results in false negative, as in (a).
+In contrast, our TRD employs a multi-branch design and Crossmodal Tuners to preserve modality-specific anomaly cues, enabling more accurate anomaly localization as in (b).
+Currently, feature learning-based methods for multimodal AD are primarily divided into two categories.
+One is Crossmodal Feature Mapping [6], which introduces a novel paradigm for addressing multimodal AD, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships.
+When anomalies occur, the failure of feature mapping enables anomaly localization.
 But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
+Knowledge distillation-based methods [5, 9] represent another widely accepted category of feature learning-based methods, which extend the typical KD paradigm, originally used in image anomaly detection, to multimodal settings.
+In such methods, student networks are trained on normal samples to learn multimodal feature representations from teacher networks.
+Anomalies are then identified by measuring the feature discrepancies between the student and the teacher.
+To obtain multimodal teacher features as distillation targets, KD-based multimodal AD methods typically adopt one of two fusion strategies.
+Some fuse 2D and 3D data at the input level before passing it through a pre-trained encoder.
+Others extract features from separately frozen 2D and 3D encoders and then merge them.
+However, both fusion strategies smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies, particularly when one modality is normal and the other is anomalous.
+In such cases, the fused teacher features tend to be dominated by the normal modality, reducing their sensitivity to subtle anomalies and increasing the risk of false negatives, as illustrated in Fig.
+2 (a).
 To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one.
-Each modality has its own distillation target, ensuring the sensitivity to anomalies within each modality.
-Additionally, to tackle the underfitting problem of Crossmodal Feature Mapping, we propose not to directly use the feature similarity from crossmodal mapping for anomaly map generation.
-Instead, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
-In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [13], named Crossmodal Reverse Distillation (CRD).
-CRD naturally combines the strengths of both feature learning paradigms.
+Each modality is assigned an independent distillation objective, allowing the student network to learn from the corresponding teacher features.
+This multi-branch design preserves the anomaly sensitivity of each modality and avoids the anomaly smoothing problem caused by early feature fusion.
+With this setup, each branch is able to accurately detect anomalies specific to its modality.
+However, simply performing late fusion by combining anomaly maps from each branch may still weaken anomaly signals, particularly when anomalies only in one modality are subtle.
+To further mitigate this issue, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
+This enables each branch not only to retain its own anomaly information but also to capture crossmodal cues, ensuring that anomalies present in any modality are effectively reflected in the final output.
+In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [10], named Tuned Reverse Distillation (TRD).
+TRD naturally combines the strengths of existing feature learning paradigms, as shown in Fig.
+2 (b).
 First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps.
-Next, we introduce two crossmodal assistants within each modality branch.
+Next, we introduce two Crossmodal Tuners within each modality branch.
 One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features.
@@
 Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches.
-Experimental results on representative multimodal AD datasets demonstrate the effectiveness of CRD in multimodal anomaly detection and localization.
-2 Related Work
-Unsupervised Anomaly Detection.
+Experimental results on representative multimodal AD datasets demonstrate the effectiveness of TRD in multimodal anomaly detection and localization.
+Our contributions are as follows:
+  • We propose Tuned Reverse Distillation based on Multi-branch Distillation for multimodal industrial AD, which effectively detects anomalies in all modalities.
+  • Two Crossmodal Tuners are introduced, including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
+  • Experiments show that our proposed TRD achieves state-of-the-art performance on multimodal AD datasets.
+II Related Work
+II-A Unsupervised Anomaly Detection
 Unsupervised AD has gained increasing attention in recent years.
-Initially, many unsupervised AD methods relied on generative models [5, 1, 38, 32, 19, 18].
+Initially, many unsupervised AD methods relied on generative models [11, 12, 13, 14, 15, 16, 17].
 These models are trained on normal samples to learn the ability to reconstruct normal data.
@@
 Recently, significant progress has been made in the research of synthetic anomaly images [21, 22, 23], which simulate real-world scenarios to assist in unsupervised AD tasks.
-Additionally, knowledge distillation methods [27, 30, 26, 3, 13, 28, 16, 14, 37, 17] based on the teacher-student framework have also been applied to unsupervised AD.
+Additionally, knowledge distillation methods [24, 5, 10, 25, 26, 27, 28, 29] based on the teacher-student framework have also been applied to unsupervised AD.
 These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies.
 Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
-Multimodal Anomaly Detection.
+II-B Multimodal Anomaly Detection
 With the release of multiple 3D industrial anomaly detection datasets [4, 30, 31], unsupervised multimodal AD has gradually become a topic of research.
@@
 Crossmodal Feature Mapping [6] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
-3 Preliminaries
+III Preliminaries
 Knowledge distillation (KD) is a widely recognized and researched paradigm for unsupervised industrial image AD, typically based on a teacher-student network framework.
@@
 Hence, even for anomalous samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
-4 Crossmodal Reverse Distillation
-Figure 3: Pipeline of our proposed method.
-The feature evolution visualization highlights that anomalies in each modality are effectively captured.
-In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
-Figure 4: Information flow of the branch of Modality A, showing how Crossmodal Filter and Crossmodal Amplifier incorporate information from Modality B.
-Crossmodal Filter aids the student decoder in generating normal features by utilizing features from Modality B, while Crossmodal Amplifier strengthens the perception of anomalies in Modality A by amplifying anomaly signals from Modality B.
+IV Tuned Distillation
+Figure 3: Information flow of the branch for Modality A during training.
+This branch consists of a frozen teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier.
+Crossmodal Filter incorporates mapped features from Modality B to guide the student decoder of Modality A in generating anomaly-free features.
+Crossmodal Amplifier injects anomaly signals from Modality B after the decoder, enhancing the anomaly sensitivity of Modality A.
 In the multimodal AD task, each input consists of multiple modalities of data.
@@
 The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$.
-The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
-4.1 Multi-branch Distillation
-In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories: (1) Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem [26].
-(2) Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network [15].
+The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which include both normal and anomalous samples.
+IV-A Multi-branch Distillation
+In multimodal industrial AD, several methods have incorporated Knowledge Distillation, which are broadly classified into two categories:
+  • Some methods add or concatenate multimodal data at the input stage, transforming the multimodal into a singlemodal problem [5].
+  • Others employ multiple teacher networks to independently extract features from different modalities, and then merge the features as the learning target for the distillation of the student network [9].
 The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern.
-During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
+During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features of the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
 However, the effectiveness of these methods relies on a key assumption: the fused teacher features exhibit strong sensitivity to anomalies, which means they generate a significant difference between normal and anomalous regions.
@@
 To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig.
-3, which differs from previous KD-based multimodal AD methods with a multi-branch design.
+2 (b), which differs from previous KD-based multimodal AD methods with a multi-branch design.
 (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality.
@@
 For the design of each branch, we draw inspiration from RD.
-Both 2D and 3D branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec.
-4.2).
-The teacher encoders in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet.
-The student decoders follow RD, using a structure symmetric to that of the teacher networks.
-Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to optimize the output features, enabling crossmodal mapping and information fusion between modalities.
-The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
-4.2 Crossmodal Assistants
+The two branches corresponding to the 2D and 3D modalities are designed symmetrically.
+Fig.
+3 depicts the detailed architecture of one branch (for modality A), which includes four components: a frozen pre-trained teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier (Section IV-B).
+The teacher encoder in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet.
+The student decoders in both branches follow RD, using a structure symmetric to that of the teacher networks.
+Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to tune the output features, enabling crossmodal mapping and information fusion between modalities.
+Figure 4: Illustration of the Crossmodal Filter within the branch of Modality A, which consists of two components including Bottleneck Projection and OCBE.
+First, the encoder features from the other modality (Modality B) ${F_{E}^{B}}^{i},i=\{1,2,3\}$ are processed by the Bottleneck Projection to suppress anomaly signals and generate compact, normal-like features.
+These features are then concatenated with the encoder features from Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ and passed into the OCBE module.
+The resulting fused representation $F_{CF}^{A}$ retains rich multimodal information while remaining anomaly-free, and is used to guide the student decoder in reconstructing anomaly-free features within the Modality A branch.
+IV-B Crossmodal Tuners
 To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation.
-To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies.
+To be specific, we design two Crossmodal Tuners serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies.
 (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality.
 With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
-Crossmodal Filter.
+IV-B1 Crossmodal Filter
 The training objective of RD is to ensure that the student decoder generates anomaly-free features.
@@
 It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration.
-First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $8\times 8$ by convolution.
+First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $\times 8$ 8 by convolution.
 Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$.
@@
 During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
-Crossmodal Amplifier.
-When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec.
-4.1.
+Figure 5: Illustration of the Crossmodal Amplifier within the branch of Modality A.
+The module is primarily composed of an Inverted Bottleneck Projection.
+Encoder features ${F_{E}^{B}}^{i},i=\{1,2,3\}$ from the other modality (Modality B) are first processed by this module to generate anomaly-aware features.
+These features are designed to resemble the encoder features of Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ in normal regions of Modality B, while exhibiting discrepancies in anomalous regions.
+This contrast enables the anomaly information from Modality B to be reflected in the Modality A branch.
+Finally, the amplified features are added to the student decoder output of Modality A to produce the final outputs ${F_{CA}^{B}}^{i},i=\{1,2,3\}$.
+IV-B2 Crossmodal Amplifier
+When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Section IV-A.
 We address this issue with Multi-branch Distillation.
@@
 As illustrated in Fig.
-4, we design Crossmodal Amplifier.
-First, Inverted Bottleneck Projection (IBP) is designed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$.
-Then, the mapped features are then dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by
+5, we design Crossmodal Amplifier.
+First, Inverted Bottleneck Projection (IBP) is proposed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$.
+Then, the mapped features are dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by
 ${F^{A}_{CA}}^{i}=\frac{\exp({w_{1}^{A}}^{i})*{F^{A}_{D}}^{i}+\exp({w_{2}^{A}}^{i})*{F^{B}_{IBP}}^{i}}{\exp({w_{1}^{A}}^{i})+\exp({w_{2}^{A}}^{i})}$ (3)
@@
 This process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information during feature mapping.
-Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Avg.
-DepthGAN [7] 53.8/42.1 37.2/42.2 58.0/77.8 60.3/69.6 43.0/49.4 53.4/25.2 64.2/28.5 60.1/36.2 44.3/40.2 57.7/63.1 53.2/47.4
-DepthAE [7] 64.8/43.2 50.2/15.8 65.0/80.8 48.8/49.1 80.5/84.1 52.2/40.6 71.2/26.2 52.9/21.6 54.0/71.6 55.2/47.8 59.5/48.1
-DepthVM [7] 51.3/38.8 55.1/32.1 47.7/19.4 58.1/57.0 61.7/40.8 71.6/28.2 45.0/24.4 42.1/34.9 59.8/26.8 62.3/33.1 55.5/33.5
-BTF [20] 91.8/97.6 74.8/96.9 96.7/97.9 88.3/97.3 93.2/93.3 58.2/88.8 89.6/97.5 91.2/98.1 92.1/95.0 88.6/97.1 86.5/95.9
-M3DM [31] 99.4/97.0 90.9/97.1 97.2/97.9 97.6/95.0 96.0/94.1 94.2/93.2 97.3/97.7 89.9/97.1 97.2/97.1 85.0/97.5 94.5/96.4
-Shape-guided [11] 98.6/98.1 89.4/97.3 98.3/98.2 99.1/97.1 97.6/96.2 85.7/97.8 99.0/98.1 96.5/98.3 96.0/97.4 86.9/97.5 94.7/97.6
-AST [26] 98.3/97.0 87.3/94.7 97.6/98.1 97.1/93.9 93.2/91.3 88.5/90.6 97.4/97.9 98.1/98.2 100/88.9 79.7/94.0 93.7/94.4
-MMRD [15] 99.9/98.6 94.3/99.0 96.4/99.1 94.3/95.1 99.2/99.0 91.2/90.1 94.9/99.0 90.1/99.0 99.4/98.7 90.1/98.2 95.0/97.6
-CFM [12] 98.8/98.0 87.5/96.6 98.4/98.2 99.2/94.7 99.7/95.9 92.4/96.7 96.4/98.2 94.9/98.3 97.9/97.6 95.0/98.2 96.0/97.2
-CRD (Ours) 99.6/97.7 96.1/98.1 98.2/99.1 98.5/98.2 99.7/98.7 90.1/91.0 99.2/99.4 97.6/99.5 100/98.2 83.3/97.4 96.2/97.7
-Table 1: Anomaly detection and localization results I-AUROC/PRO (%) on MVTec 3D-AD with the best in bold and the second-best underlined.
-4.3 Training Objectives
+IV-C Training Objectives
 The overall optimization objective of our method is composed of three main components.
@@
 where $\mathcal{L}^{2D}_{D}$ and $\mathcal{L}^{3D}_{D}$ are the losses corresponding to 2D and 3D branches respectively.
-Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ within these filters to the encoder features.
-The training losses of 2D and 3D branches are calculated as:
+Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ output by Crossmodal Filters to the encoder features, as shown in Fig.
+4.
+The training losses of 2D and 3D branches are calculated as
 $\mathcal{L}^{2D/3D}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{BP}^{3D/2D}}^{i})\}$ (5)
-Finally, Crossmodal Amplifier needs to be optimized,, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. The total loss function is
+Finally, Crossmodal Amplifier needs to be optimized, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$.
+To guide this process, as in Fig.
+5, we define a total loss for Crossmodal Amplifier composed of two parts: the feature mapping loss $\mathcal{L}^{2D/3D}_{IBP}$, which aligns the mapped crossmodal features with the target encoder features, and the output consistency loss $\mathcal{L}^{2D/3D}_{output}$, which is used to adjust feature fusion parameters.
+The total loss function is
 $\mathcal{L}^{2D/3D}_{CA}=\mathcal{L}^{2D/3D}_{IBP}+\mathcal{L}^{2D/3D}_{output}$ (6)
@@
 $\mathcal{L}^{2D/3D}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i})\}$ (8)
-The overall loss $\mathcal{L}_{CRD}$ is expressed as
-$\mathcal{L}_{CRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
-Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.
-4.4 Anomaly Scoring
+The overall loss $\mathcal{L}_{TRD}$ is expressed as
+$\mathcal{L}_{TRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
+Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.0.
+Figure 6: Inference process of our proposed method.
+The feature evolution visualization highlights that anomalies in each modality are effectively captured.
+In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
+TABLE I: Anomaly Detection and Localization Results (%) on MVTec 3D-AD with the Best in Bold and the Second-best Underlined
+Metrics Method Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
+I-AUC BTF [7] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
+M3DM [8] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
+Shape-guided [32] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
+AST [5] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
+MMRD [9] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
+CFM [6] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
+TRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
+P-AUC BTF [7] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
+M3DM [8] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
+AST [5] - - - - - - - - - - 97.6
+MMRD [9] - - - - - - - - - - 99.2
+CFM [6] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
+TRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
+PRO BTF [7] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
+M3DM [8] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
+Shape-guided [32] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
+AST [5] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
+MMRD [9] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
+CFM [6] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
+TRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
+IV-D Inference
+IV-D1 Anomaly Scoring
 Following previous KD-based methods, we generate the anomaly map for each branch using cosine distance as
@@
 And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
-Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
-M3DM [31] 62.4/90.6 95.8/92.3 95.8/80.3 100/98.3 88.6/85.5 75.8/68.8 94.9/88.0 83.6/90.6 100/96.6 100/95.5 89.7/88.2
-AST [26] 57.4/51.4 74.7/83.5 74.7/71.4 88.9/90.5 59.6/58.7 61.7/59.0 81.6/73.6 84.1/76.9 98.7/91.8 98.7/87.8 78.0/74.4
-MMRD [15] 85.4/97.5 100/97.0 94.6/94.2 99.8/98.5 90.8/91.7 74.7/68.0 96.6/97.0 98.4/94.1 100/99.0 100/99.2 94.0/93.6
-CFM [12] 68.0/94.2 93.1/90.2 95.2/83.1 88.0/96.5 86.5/87.5 78.2/76.2 91.7/79.1 84.0/91.3 99.8/93.9 96.2/94.9 88.1/88.7
-CRD (Ours) 86.7/98.0 100/94.1 98.1/96.8 98.1/98.1 87.2/93.6 72.8/77.3 97.1/98.1 93.0/94.4 100/98.9 99.5/98.1 93.3/94.7
-Table 2: Anomaly detection and localization results I-AUC/PRO (%) on Eyecandies with the best in bold and the second-best underlined.
-5 Experiments
-5.1 Experimental Settings
-Dataset.
-We conduct experiments MVTec 3D-AD [7] and Eyecandies [9].
+IV-D2 Feature Evolution Analysis
+To better understand the behavior of our TRD during inference, Fig.
+6 visualizes the feature evolution process within TRD.
+Intuitively, Crossmodal Filter effectively supplements information from the other modality while suppressing anomaly signals that may interfere with anomaly-free feature reconstruction.
+In contrast, Crossmodal Amplifier allows anomaly information from one modality to be reflected in the other branch.
+Through the joint modulation of both components, each branch becomes aware of all modality-specific anomalies, resulting in more accurate and robust anomaly detection.
+V Experiments
+V-A Experimental Settings
+Dataset
+We conduct experiments on datasets MVTec 3D-AD [4] and Eyecandies [30].
 MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects.
@@
 In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
-Evaluation Metrics.
+Evaluation Metrics
 The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP).
@@
 We also employ the per-region overlap (PRO) metric calculated according to the official code [4] for anomaly localization, which treats anomaly regions of different size equally.
-Implementation Details.
-All RGB images, depth maps and normal maps are scaled to $256\times 256$.
+Implementation Details
+All RGB images, depth maps and normal maps are scaled to $\times 256$ 256.
 The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category.
@@
 In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
-Figure 5: Qualitative ablation results of crossmodal assistants on MVTec 3D-AD.
-5.2 Main Results
-Results on MVTec 3D-AD.
-Tab.
-1 shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset.
-We compare our method CRD with several state-of-the-art 3D anomaly detection methods.
-CRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%.
-Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach.
+TABLE II: Anomaly Detection and Localization Results (%) on Eyecandies with the Best in Bold and the Second-best Underlined
+Metrics Methods
+Candy
+Cane
+Chocolate
+Cookie
+Chocolate
+Praline
+Confetto
+Gummy
+Bear
+Hazelnut
+Truffle
+Licorice
+Sandwich
+Lollipop
+Marsh-
+mallow
+Peppermint
+Candy
+Average
+I-AUC M3DM [8] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
+AST [5] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
+MMRD [9] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
+CFM [6] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
+TRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
+P-AUC M3DM [8] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
+AST [5] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
+MMRD [9] - - - - - - - - - - 98.3
+CFM [6] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
+TRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
+PRO M3DM [8] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
+AST [5] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
+MMRD [9] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
+CFM [6] 94.2 90.2 83.1 88.0/96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
+TRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
+V-B Main Results
+V-B1 Results on MVTec-3D AD
+Table I shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset.
+We compare our method TRD with several state-of-the-art 3D anomaly detection methods.
+TRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%.
+Notably, TRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach.
 A subset of qualitative results on MVTec 3D-AD is provided in Fig.
-1, showing that our proposed CRD accurately localizes the anomaly regions.
-Results on Eyecandies.
-We evaluate our method on the Eyecandies dataset, as shown in Tab.
-2.
-Our method CRD significantly outperforms all non-KD-based multimodal anomaly detection methods.
-While it slightly lags behind MMRD [15] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
-Method MVTec 3D-AD (%) Eyecandies (%) Avg. (%) Frame Rate (fps)
+1, showing that our proposed TRD accurately localizes the anomaly regions.
+Figure 7: Performance comparison of feature learning-based methods via radar chart.
+TABLE III: Comparison of Feature Learning-based Methods on MVTec 3D-AD and Eyecandies datasets under the Same Hardware Setup (* indicates that MMRD has no reproducible code)
+Method MVTec 3D-AD (%) Eyecandies (%)
+Avg.
+Frame
+Rate (fps)
 I-AUC P-AUC PRO I-AUC P-AUC PRO
@@
 CFM [6] 96.0 99.3 97.2 88.1 97.4 88.7 94.5 12.0
-CRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
-Table 3: Comparison of feature learning-based methods on MVTec 3D-AD and Eyecandies datasets under the same hardware setup. * indicates that MMRD has no reproducible code.
-Comparison with Feature Learning-based Methods.
-We further compare CRD with feature learning-based methods including AST[26], MMRD[15], and CFM [12] on MVTec 3D-AD and Eyecandies.
-As in Table 3, CRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps.
-Despite the multi-branch design, CRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
-5.3 Ablation Studies
+TRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
+V-B2 Results on Eyecandies
+We evaluate our method on the Eyecandies dataset, as shown in Table II.
+Our method TRD significantly outperforms all non-KD-based multimodal anomaly detection methods.
+While our TRD slightly lags behind MMRD [9] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
+V-B3 Comparison with Feature Learning-based Methods
+We further compare TRD with feature learning-based methods including AST[5], MMRD[9], and CFM [6] on MVTec 3D-AD and Eyecandies.
+As in Fig.
+7 and Table 7, TRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps.
+Despite the multi-branch design, TRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
+TABLE IV: Anomaly Detection and Localization Results (%) of Different Modality Branches on MVTec 3D-AD
 Branch I-AUC P-AUC P-PRO I-AP P-AP Avg.
 2D 86.8 98.9 95.4 95.8 37.7 82.9
-2D+Crossmodal 88.9 99.0 95.8 96.5 39.4 83.9
+2D+CT 88.9 99.0 95.8 96.5 39.4 83.9
 3D 87.1 96.2 88.8 95.2 33.7 80.2
-3D+Crossmodal 87.3 97.1 90.1 95.7 34.1 80.9
+3D+CT 87.3 97.1 90.1 95.7 34.1 80.9
 2D+3D 93.9 99.3 97.4 98.0 45.2 86.8
-2D+3D+Crossmodal 96.2 99.4 97.7 98.8 45.5 87.5
-Table 4: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
+2D+3D+CT (Ours) 96.2 99.4 97.7 98.8 45.5 87.5
+Figure 8: Visualization of anomaly maps of different modality branches on MVTec 3D-AD.
+Figure 9: Visualization of anomaly maps of different modality branches on Eyecandies.
+V-B4 Singlemodal vs. Multimodal
+We investigate the effectiveness of combining different modalities for AD through both quantitative and qualitative comparisons.
+We first compare singlemodal baselines (2D and 3D), their enhanced versions with Crossmodal Tuners (2D+CT and 3D+CT), and the full multimodal settings (2D+3D and 2D+3D+CT) in Table IV.
+Results on the MVTec 3D-AD dataset show that combining both modalities significantly improves performance.
+The I-AUC metric of 93.9% for using multiple modalities is higher than 86.8% for 2D only and 87.1% for 3D only, highlighting the complementary benefits of modality fusion.
+To provide further insight, we visualize the anomaly maps of each branch under multimodal setup on MVTec 3D-AD and Eyecandies, as in Fig.
+8 and Fig.
+9.
+Some anomalies are hard to perceive in one modality but clearly visible in the other.
+Relying on a single modality may lead to missed detections, while our TRD designed for multimodal AD helps mitigate this issue through complementary information across modalities.
+V-C Ablation Studies
+TABLE V: Ablation Study Results (%) of Crossmodal Tuners on MVTec 3D-AD
 CF CA I-AUC P-AUC P-PRO I-AP P-AP Avg.
@@
 96.2 99.4 97.7 98.8 45.5 87.5
-Table 5: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
-Study on Multi-branch Distillation.
-We conduct an ablation study in Tab.
-4, which demonstrates that the multi-branch setup, combining both 2D and 3D modalities, significantly outperforms individual modality branches.
-The I-AUC metric of 93.9% for using both branches is higher than 86.8% for 2D branch only and 87.1% for 3D branch only, highlighting the complementary benefits of modality fusion.
-Study on Crossmodal Assistants.
-We perform an ablation study to evaluate the impact of crossmodal assistants on anomaly detection performance, as shown in Tab.
-5.
+Figure 10: Qualitative ablation results of Crossmodal Tuners on MVTec 3D-AD.
+V-C1 Study on Crossmodal Tuners
+We perform an ablation study to evaluate the impact of Crossmodal Tuners on anomaly detection performance, as shown in Table V.
 The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results.
@@
 The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
-We also provide visualizations of anomaly maps before and after incorporating crossmodal assistants, as in Fig.
-5.
-Without crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch.
-However, after introducing crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
+We also provide visualizations of anomaly maps before and after incorporating Crossmodal Tuners, as in Fig.
+10.
+Without Crossmodal Tuners, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch.
+However, after introducing Crossmodal Tuners, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
+TABLE VI: Ablation Study Results (%) of Different Downsampling Sizes in Bottlneck Projection on MVTec 3D-AD
 Downsampling Size I-AUC P-AUC P-PRO I-AP P-AP Avg.
 Identity 93.9 99.3 97.4 98.0 45.8 86.9
-$16\times 16$ 93.6 99.3 97.4 97.9 45.2 86.7
-$8\times 8$ 94.2 99.3 97.5 98.2 45.8 87.0
-$4\times 4$ 93.8 99.3 97.4 98.0 45.3 86.8
-Table 6: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
-Study on Bottlneck Projection in Crossmodal Filter.
-We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab.
-6.
-When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics.
-It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively.
+$\times 16$ 16 93.6 99.3 97.4 97.9 45.2 86.7
+$\times 8$ 8 94.2 99.3 97.5 98.2 45.8 87.0
+$\times 4$ 4 93.8 99.3 97.4 98.0 45.3 86.8
+V-C2 Study on Bottlneck Projection in Crossmodal Filter
+We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Table VI.
+When Bottleneck Projection reduces the feature size to $\times 8$ 8, the model achieves the best performance across multiple metrics.
+It is observed that excessively large downsampling size (e.g., $\times 4$ 4) leads to information loss, while small downsampling size (e.g., $\times 16$ 16) makes it difficult to filter out abnormal interference effectively.
 Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
+TABLE VII: Ablation Study Results (%) of Different Channel Expansion Sizes in Inverted Bottlneck Projection on MVTec 3D-AD
 Channel Expansion I-AUC P-AUC P-PRO I-AP P-AP Avg.
@@
 4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0 87.3
-Table 7: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
-Study on Inverted Bottlneck Projection in Crossmodal Amplifier.
-As recorded in Tab.
-7, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier.
+V-C3 Study on Inverted Bottlneck Projection in Crossmodal Amplifier
+As recorded in Table VII, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier.
 The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency.
 This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
+TABLE VIII: Ablation Study Results (%) of Anomaly Map Aggregation Strategy on MVTec 3D-AD
 Aggregation I-AUC P-AUC P-PRO I-AP P-AP Avg.
@@
 $\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [6] 94.3 99.3 97.2 98.2 43.5 86.5
-Table 8: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
-Study on Anomaly Scoring.
-Tab.
-8 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps.
+V-C4 Study on Anomaly Scoring
+Table VIII demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps.
 Normalization ensures that anomaly maps from different modalities are on the same scale, making fusion more effective.
 Figure 11: Visualization of the features output by CF and CA.
-Visualization of Mapped Features.
+V-C5 Visualization of Mapped Features
 We visualizes the output features of CF and CA using the visualization decoders [13] in Fig.
-6, which intuitively demonstrates the roles of CA and CF.
+11, which intuitively demonstrates the roles of CA and CF.
 Regardless of whether the initial modality is normal or anomalous, CF consistently produces normal outputs, enabling the student decoder of the other modality to generate anomaly-free results.
 When the initial modality is anomalous, CA introduces discrepancies with the other modality, amplifying anomalies for enhanced detection.
-6 Conclusion
-In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task.
-CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch.
+VI Conclusion
+In this paper, Tuned Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task.
+TRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch.
 Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities.
-Experimental results show that CRD achieves SOTA performance in both anomaly detection and localization.
-Supplementary Material
-Overview
-This supplementary material provides extended analyses and visualizations to complement the main paper:
-  • Detailed P-AUC results for anomaly localization on MVTec 3D-AD and Eyecandies datasets. (Appendix A)
-  • Training accuracy curves to demonstrate the model’s convergence over training. (Appendix B)
-  • Detailed results of ablation studies on MVTec 3D-AD dataset. (Appendix C)
-  • Additional qualitative visualizations on MVTec 3D-AD and Eyecandies datasets. (Appendix D)
-Appendix A P-AUC Metric
-A.1 P-AUC on MVTec 3D-AD
-Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-BTF [20] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
-M3DM [31] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
-AST [26] - - - - - - - - - - 97.6
-MMRD [15] - - - - - - - - - - 99.2
-CFM [12] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
-CRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-Table 9: Pixel-level anomaly localization results P-AUC (%) on MVTec 3D-AD with the best in bold.
-This section compares our method with other mainstream multimodal AD methods on MVTec 3D-AD dataset using the P-AUC metric, including memory bank-based methods BTF [20] and M3DM [31], KD-based methods AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
-As shown in Tab.
-9, our method achieves SOTA performance, reaching 99.4% on the P-AUC metric.
-Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods.
-In most other categories, it consistently maintains competitive or near-optimal performance.
-A.2 P-AUC on Eyecandies
-Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
-RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
-RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
-M3DM [31] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
-AST [26] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
-MMRD [15] - - - - - - - - - - 98.3
-CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
-CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
-Table 10: Pixel-level anomaly localization results P-AUC (%) on Eyecandies with the best in bold.
-This section presents the quantitative results of anomaly localization on Eyecandies dataset using the P-AUC metric in Tab.
-10.
-The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [20] and M3DM [31], KD-based methods like AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12].
-Our method significantly outperforms all non-KD-based multimodal anomaly detection methods, achieving 98.5%.
-Appendix B Training Curves: I-AUC, P-AUC, and PRO
-(a) I-AUC.
-(b) P-AUC.
-(c) PRO.
-Figure 7: AD performance over 500 training epochs.
-We provide training curves of I-AUC, P-AUC, and PRO for the carrot and peach categories from the MVTec 3D-AD dataset.
-As shown in Fig.
-7, all metrics stabilize after approximately 50 epochs, indicating that the model converges efficiently without significant performance fluctuations.
-This case demonstrates the robustness of our method, as it avoids over-generalization even with extended training.
-Following the setting of Reverse Distillation [13], we adopt 200 epochs for training to ensure sufficient convergence while maintaining computational efficiency.
-Appendix C Additional Results of Ablation Studies
-In the main text, we present five aspects of ablation studies:
-  1. 1.
-    Study on Multi-branch Distillation demonstrates that combining the distillation results from multiple branches of different modalities achieves better performance, highlighting the importance of multiple branches.
-(Tab.
-11)
-  2. 2.
-    Study on Crossmodal Assistants shows that incorporating crossmodal assistants into Multi-branch Distillation based on Reverse Distillation significantly improves anomaly detection, proving their effectiveness.
-(Tab.
-12)
-  3. 3.
-    Study on Bottleneck Projection in Crossmodal Filter explores the impact of downsampling sizes in the Bottleneck Projection module of Crossmodal Filter, concluding that a size of $8\times 8$ yields optimal results.
-(Tab.
-13)
-  4. 4.
-    Study on Inverted Bottleneck Projection in Crossmodal Amplifier investigates the effect of channel expansion in the Inverted Bottleneck Projection module of Crossmodal Amplifier, finding that a 2×2$\times$2 × expansion achieves the best performance.
-(Tab.
-14)
-  5. 5.
-    Study on Anomaly Scoring examines different methods for fusing anomaly maps from multiple branches, confirming that our selected approach delivers the best results. (Tab. 15)
-In the main text, we provide average results across all categories for comparison. Here, we present detailed results for each category, as shown in Tab. 11, Tab. 12, Tab. 13, Tab. 14, and Tab. 15.
-Metrics Branch Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUC 2D 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
-2D+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
-3D 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
-3D+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
-2D+3D 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
-2D+3D+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-P-AUC 2D 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
-2D+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
-3D 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
-3D+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
-2D+3D 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
-2D+3D+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-PRO 2D 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
-2D+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
-3D 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
-3D+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
-2D+3D 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
-2D+3D+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
-I-AP 2D 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
-2D+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
-3D 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
-3D+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
-2D+3D 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
-2D+3D+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
-P-AP 2D 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
-2D+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
-3D 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
-3D+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
-2D+3D 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
-2D+3D+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
-Table 11: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
-Metrics CF CA Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUC - - 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
-- 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
-- 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
-99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-P-AUC - - 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
-- 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
-- 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
-99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-PRO - - 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
-- 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
-- 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
-97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
-I-AP - - 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
-- 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
-- 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
-99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
-P-AP - - 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
-- 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
-- 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
-27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
-Table 12: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
-Metrics Downsampling Size Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUC Identity 99.6 88.8 97.9 95.6 99.7 84.6 98.8 96.8 99.8 77.1 93.9
-$16\times 16$ 99.7 88.1 97.6 96.9 100 81.6 99.2 96.7 99.8 76.7 93.6
-$8\times 8$ 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
-$4\times 4$ 99.7 89.5 98.1 95.6 99.9 81.6 98.8 96.8 99.8 77.5 93.8
-P-AUC Identity 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.1 99.3
-$16\times 16$ 99.4 99.3 99.7 99.3 99.7 97.3 99.8 99.9 99.6 99.1 99.3
-$8\times 8$ 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
-$4\times 4$ 99.4 99.3 99.7 99.3 99.6 97.0 99.8 99.9 99.6 99.1 99.3
-PRO Identity 98.2 98.0 99.0 98.1 98.4 90.2 99.3 99.5 98.0 95.6 97.4
-$16\times 16$ 98.1 98.0 99.0 98.1 98.4 90.6 99.2 99.5 98.0 95.5 97.4
-$8\times 8$ 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
-$4\times 4$ 98.1 98.1 99.0 98.0 98.4 89.8 99.2 99.5 98.0 95.7 97.4
-I-AP Identity 99.9 96.9 99.5 98.8 99.9 95.5 99.7 99.0 99.9 91.3 98.0
-$16\times 16$ 99.9 96.7 99.5 99.2 100 93.9 99.8 98.9 99.9 90.9 97.9
-$8\times 8$ 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
-$4\times 4$ 99.9 97.3 99.6 98.8 100 93.8 99.7 98.9 99.9 91.8 98.0
-P-AP Identity 45.7 37.5 41.2 69.3 40.6 27.8 62.2 57.5 52.8 23.2 45.8
-$16\times 16$ 42.7 37.1 41.4 70.2 41.3 27.3 59.6 56.8 52.9 22.3 45.2
-$8\times 8$ 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
-$4\times 4$ 44.4 38.0 41.6 70.1 40.6 26.4 59.3 56.1 54.0 22.8 45.3
-Table 13: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
-Metrics Channel Expansion Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUC Identity 99.6 95.2 98.3 98.0 99.7 87.4 99.2 97.6 100 81.9 95.7
-2×2$\times$2 × 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
-4×4$\times$4 × 99.5 94.7 98.0 98.6 99.7 88.8 99.2 97.5 100 80.8 95.7
-P-AUC Identity 99.3 99.4 99.7 99.3 99.7 97.4 99.8 99.5 99.7 99.4 99.3
-2×2$\times$2 × 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
-4×4$\times$4 × 99.3 99.3 99.7 99.3 99.7 97.6 99.8 99.9 99.7 99.4 99.4
-PRO Identity 97.6 98.3 99.1 98.2 98.6 90.9 99.3 99.5 98.2 97.0 97.7
-2×2$\times$2 × 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
-4×4$\times$4 × 97.7 98.2 99.1 98.2 98.6 91.6 99.3 99.5 98.2 96.8 97.7
-I-AP Identity 99.9 98.9 99.6 99.5 99.9 96.5 99.8 99.3 100 93.2 98.7
-2×2$\times$2 × 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
-4×4$\times$4 × 99.9 98.7 99.5 99.6 99.9 96.9 99.8 99.3 100 92.8 98.6
-P-AP Identity 27.9 39.9 42.1 67.6 49.8 28.4 62.0 57.0 55.4 21.3 45.1
-2×2$\times$2 × 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
-4×4$\times$4 × 29.0 38.4 42.4 68.1 48.8 29.1 60.9 57.3 55.7 20.4 45.0
-Table 14: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
-Metrics Aggregation Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
-I-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.6 92.9 98.3 98.3 99.3 90.3 98.9 96.7 100 81.6 95.6
-$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
-P-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.4 96.8 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.1
-$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
-PRO $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 97.8 96.8 99.1 98.2 98.7 91.1 99.3 99.5 98.2 97.2 97.6
-$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
-I-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.9 98.2 99.6 99.6 99.8 97.2 99.7 99.0 100 92.7 98.6
-$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
-P-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
-$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 31.6 27.0 43.9 70.4 48.8 29.6 60.3 55.9 55.3 21.2 44.4
-$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
-Table 15: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
-Appendix D Additional Qualitative Results
-Figure 8: Qualitative results on MVTec 3D-AD.
-From top to bottom: 2D images, 3D depth maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
-D.1 Qualitative Results on MVTec 3D-AD
-This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig.
-8.
-The visualizations reveal two key findings: (1) In the 3D branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives.
-But this issue is less pronounced in the 2D branch.
-(2) Some anomalies are modality-specific.
-For instance, discoloration is only detectable in RGB images, while the absence of chocolate on cookies is only evident in depth maps.
-Consequently, anomaly maps from individual branches may fail to fully capture all anomalies within a sample.
-However, the fused results effectively achieve comprehensive anomaly localization.
-Figure 9: Qualitative results on MVTec 3D-AD. From left to right: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
-In the main text, we visualize anomaly maps with and without crossmodal assistants, demonstrating their effectiveness in improving anomaly segmentation accuracy by reducing false positives and false negatives.
-In this section, we provide additional examples to further validate the robustness of our method.
-As shown in Fig.
-9, the inclusion of CF and CA consistently enhances the precision of anomaly localization across various categories, including bagel, cable cland, cookie, dowel, and potato.
-These results highlight the critical role of crossmodal interactions in refining anomaly detection.
-D.2 Qualitative result on Eyecandies
-Figure 10: Qualitative results on Eyecandies.
-From top to bottom: the input 2D images, the corresponding 3D normal maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
-Fig.
-10 visualizes the anomaly maps generated by our proposed CRD on Eyecandies.
-Our method demonstrates strong sensitivity to both 2D and 3D modalities, enabling comprehensive multimodal anomaly localization.
+Experimental results show that TRD achieves SOTA performance in both anomaly detection and localization.